### v9

# Parameter estimation for Gibbs distributions

Abstract
A central problem in computational statistics is to convert a procedure for sampling combinatorial objects into a procedure for counting those objects, and vice versa. We consider sampling problems coming from Gibbs distributions, which are families of probability distributions over a discrete space $\Omega$ with probability mass function of the form $\mu^{\Omega}_{\beta}(\omega)\propto e^{\beta H(\omega)}$ for $\beta$ in an interval $[\beta_{\min},\beta_{\max}]$ and $H(\omega)\in\{0\}\cup[1,n]$. Two important parameters are the partition function, which is the normalization factor $Z(\beta)=\sum_{\omega\in\Omega}e^{\beta H(\omega)}$, and the vector of preimage counts $c_{x}=|H^{-1}(x)|$.
We develop black-box sampling algorithms to estimate the counts using roughly $\tilde{O}(\frac{n^{2}}{\varepsilon^{2}})$ samples for integer-valued distributions and $\tilde{O}(\frac{q}{\varepsilon^{2}})$ samples for general distributions, where $q=\log\frac{Z(\beta_{\max})}{Z(\beta_{\min})}$ (ignoring some second-order terms and parameters). We show this is optimal up to logarithmic factors. We illustrate with improved algorithms for counting connected subgraphs, independent sets, and perfect matchings.
As a key subroutine, we estimate all values of the partition function using $\tilde{O}(\frac{n^{2}}{\varepsilon^{2}})$ samples for integer-valued distributions and $\tilde{O}(\frac{q}{\varepsilon^{2}})$ samples for general distributions. This improves over a prior algorithm of Huber (2015) which computes a single point estimate $Z(\beta_{\max})$ and which uses a slightly larger amount of samples. We show matching lower bounds, demonstrating this complexity is optimal as a function of $n$ and $q$ up to logarithmic terms.
This is an extended version, which includes work under the same name from ICALP 2023, as well as the earlier work [22] appearing in COLT 2018.
1 Introduction
A central problem in computational statistics is to convert a procedure for sampling combinatorial objects into a procedure for counting those objects, and vice versa. We will consider sampling algorithms for Gibbs distributions. Formally, given a real-valued function $H(\cdot)$ over a finite set $\Omega$, the Gibbs distribution is the family of distributions $\mu^{\Omega}_{\beta}$ over $\Omega$, parameterized by $\beta$ over an interval $[\beta_{\min},\beta_{\max}]$, of the form
$\mu^{\Omega}_{\beta}(\omega)=\frac{e^{\beta H(\omega)}}{Z(\beta)}$
The normalizing constant $Z(\beta)=\sum_{\omega\in\Omega}e^{\beta H(\omega)}$ is called the partition function. These distributions appear in a number of sampling algorithms; they also frequently occur in physics, where the parameter $-\beta$ corresponds to the inverse temperature, and $H(\omega)$ is called the Hamiltonian of the system.
There are two parameters of these distributions which are particularly important. The first is the vector of counts $c_{x}=|H^{-1}(x)|$ for $x\in\mathbb{R}$. This is also known as the (discrete) density of states, that is, the number of preimages of each point under $H$. In statistical physics, for instance, it essentially gives full information about the system and physically relevant quantities such as entropy, free energy, etc. A second parameter, whose role is less intuitive, is the partition ratio function
$Q({\beta})=\frac{Z(\beta)}{Z(\beta_{\min})}\qquad\qquad\text{for $\beta\in[\beta_{\min},\beta_{\max}]$}$
There is an associated probability distribution we call the gross Gibbs distribution $\mu_{\beta}(x)$ over $\mathbb{R}$ given by
$\mu_{\beta}(x)=\frac{c_{x}e^{\beta x}}{Z(\beta)},\qquad\qquad Z(\beta)=\sum_{x}c_{x}e^{\beta x}$
Note that, for the gross Gibbs distributions, the counts $c_{x}$ can only be determined up to scaling. When $c_{0}>0$, the value $\beta=-\infty$ is allowed and meaningful; in this case, we have $Z(-\infty)=c_{0}$ and $\mu_{-\infty}$ is simply the distribution which is equal to 00 with probability one. In particular, for $\beta_{\min}=-\infty$ we have $Q(\beta)=Z(\beta)/c_{0}$.
Our main aim is to develop black-box algorithms to estimate the parameters using oracle access to the gross distribution $\mu_{\beta}$ for chosen query values $\beta\in[\beta_{\min},\beta_{\max}]$. No information about the distribution $\mu_{\beta}^{\Omega}$ is used. This can simplify and unify many of the existing problem-specific estimation algorithms for Gibbs distributions across many settings.
We assume (after rescaling if necessary) that we are given known parameters $n,q$ with
$\log Q(\beta_{\max})\leq q,\qquad\qquad H(\Omega)\subseteq\mathcal{F}{\stackrel{{\scriptstyle\mbox{\tiny\tt~{}def~{}}}}{{=}}}\{0\}\cup[1,n]$
In some cases, the domain is integer-valued, i.e.
[…]
We call this the general integer setting. A special case of the integer setting, which we call the log-concave setting, is when the counts $c_{0},c_{1},c_{2},\dots,c_{n-1},c_{n}$ are non-zero and satisfy $c_{k}/c_{k-1}\geq c_{k+1}/c_{k}$ for $k=1,\dots,n-1$. The general case, where $H(\omega)$ takes unrestricted values in $\mathcal{F}$, is called the continuous setting. [^1] The log-concave algorithms still work if some of the counts $c_{i}$ are equal to zero; in this case, the non-zero counts must form a discrete interval $\{i_{0},i_{0}+1,\dots,i_{1}-1,i_{1}\}$ and the required bound must hold for $k=i_{0}+1,\dots,i_{1}-1$.
We let $\gamma$ denote the target failure probability and $\varepsilon$ the target accuracy of our algorithms, i.e. with probability at least $1-\gamma$, the algorithms should return estimates within a factor of $[e^{-\varepsilon},e^{\varepsilon}]$ of the correct value. Throughout, “sample complexity” refers to the number of calls to the oracle; for brevity, we also define the cost of an algorithm to be its expected sample complexity.
To avoid degenerate cases, we always assume $n,q\geq 2$ and $\varepsilon,\gamma\in(0,\tfrac{1}{2})$. If upper bounds $n$ and/or $q$ are not available directly, they can often be estimated by simple algorithms (up to constant factors), or can be guessed by exponential back-off strategies.
1.1 Our contributions
Our main contribution is a black-box algorithm to estimate counts. To explain this, we must address two technical issues. First, since the counts can only be recovered up to scaling in the black-box setting, some (arbitrary) normalization must be chosen. We use a convenient normalization $\pi(x)$ defined as:
[…]
The second issue is that if a count $c_{x}$ is relatively small, it is inherently hard to estimate accurately. For example, if $\mu_{\beta}(x)<\delta$ for all $\beta\in[\beta_{\min},\beta_{\max}]$ then $\Omega(1/\delta)$ samples are needed to distinguish between $c_{x}=0$ and $c_{x}>0$; with fewer samples, we will never draw $x$ from the oracle. So the complexity of estimating $c_{x}$ depends on the largest $\mu_{\beta}(x)$, over all allowed values of $\beta$. This gives rise to the parameter $\Delta(x)$ defined as
[…]
With these two provisos, we can define ${P^{\delta,\varepsilon}_{\tt count}}$ for parameters $\delta,\varepsilon\in(0,1)$ to be the problem of finding a vector $\hat{\pi}\in[0,1]^{\mathcal{F}}$ with the following two properties:
  • for all $x\in\mathcal{F}$ with $c_{x}\neq 0$, there holds $|\hat{\pi}(x)-\pi(x)|\leq\varepsilon\pi(x)(1+\frac{\delta}{\Delta(x)})$.
  • for all $x\in\mathcal{F}$ with $c_{x}=0$, there holds $\hat{\pi}(x)=0$.
In particular, if $\Delta(x)\geq\delta$, then ${P^{\delta,\varepsilon}_{\tt count}}$ provides a $(1\pm 3\varepsilon))$ relative approximation to $\pi(x)$. We emphasize here that when the algorithm succeeds (which should have probability at least $1-\gamma$), this guarantee holds simultaneously for all values $x$. We develop three main algorithmic results:
Theorem 1.
${P^{\delta,\varepsilon}_{\tt count}}$ can be solved with the following complexities:
  • In the continuous setting, with cost $O\bigl{(}\frac{\min\{q^{2}+q/\delta,q\log n+\sqrt{q\log n}/\delta\}}{\varepsilon^{2}}\log\tfrac{q}{\delta\gamma}\bigr{)}$.
  • In the general integer setting, with cost $O\bigl{(}\frac{n^{2}+n/\delta}{\varepsilon^{2}}\log^{2}\tfrac{nq}{\gamma}\bigr{)}$.
  • In the log-concave setting, with cost $O\bigl{(}\frac{\min\{(q+n)\log n,n^{2}\}+1/\delta}{\varepsilon^{2}}\log\tfrac{nq}{\gamma}\bigr{)}$.
where recall that cost refers to the expected number of queries to the oracle.
Our full results are somewhat more precise, see Theorems 2228 and 29 for more details. Note that the continuous-setting algorithm can still be applied to integer-valued problems, and is more efficient in many parameter regimes.
We also show lower bounds for black-box algorithms to solve ${P^{\delta,\varepsilon}_{\tt count}}$ by sampling the gross Gibbs distribution. We summarize these results here as follows:
Theorem 2.
Let $n\geq n_{0},q\geq q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0},\gamma<1/4$ for absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
  • $\mu$ is continuous and ${P^{\delta,\varepsilon}_{\tt count}}$ requires cost $\Omega\bigl{(}\frac{(q+\sqrt{q}/\delta)}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}$.
  • $\mu$ is integer-valued and ${P^{\delta,\varepsilon}_{\tt count}}$ requires cost $\Omega\bigl{(}\frac{\min\{q+\sqrt{q}/\delta,n^{2}+n/\delta\}}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}$.
  • $\mu$ is log-concave and ${P^{\delta,\varepsilon}_{\tt count}}$ requires cost $\Omega\bigl{(}\frac{1/\delta+\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}$.
The first two results match Theorem 1 up to logarithmic factors in $n$ and $q$. There is a small gap between the upper and lower bounds in the log-concave setting. See Theorem 47 for a more precise and general statement of these bounds.
Some theoretical count-estimation algorithms have been considered for combinatorial problems depending on specific properties of the Gibbs distribution, see e.g. [18, 9]. In physical application, the Wang-Landau (WL) algorithm [26] is a popular heuristic to estimate counts; this uses a completely different methodology from our algorithm, based on a random walk on $\mathcal{F}$ with a running count estimate $\hat{c}$. As discussed in [24], there are more than 1500 papers on the WL algorithm as well as variants such as the $1/t$-WL algorithm [3]. These algorithms are not well understood; some variants are guaranteed to converge asymptotically [10], but bounds on convergence rate or accuracy seem to be lacking. For a representative example, see for example [23], which describes a Gibbs distribution model of protein folding, and uses the WL algorithm to determine relevant properties.
Estimating the partition ratio.
As a key subroutine, we develop new algorithms to estimate partition ratios. Formally, we define ${P^{{\varepsilon}}_{\tt ratio-all}}$ to be the problem of computing a static data structure $\mathcal{D}$ with an associated deterministic function $\hat{Q}(\alpha|{\cal D})$ satisfying the property
$|\log\hat{Q}(\alpha|{\cal D})-\log Q(\alpha)|\leq\varepsilon\qquad\text{for all $\alpha\in[\beta_{\min},\beta_{\max}]$}$
We say in this case that $\mathcal{D}$ is $\varepsilon$-close. We show the following:
Theorem 3.
${P^{{\varepsilon}}_{\tt ratio-all}}$ can be solved with the following complexities:
  • In the continuous setting, with cost $O\bigl{(}\frac{\min\{q^{2},q\log n\}}{\varepsilon^{2}}\log\tfrac{1}{\gamma}\bigr{)}$.
  • In the general integer setting, with cost $O\bigl{(}\frac{n^{2}\log n}{\varepsilon^{2}}\log\tfrac{1}{\gamma}+n\log q\bigr{)}$.
  • In the log-concave integer setting, with cost $O\bigl{(}\frac{n^{2}}{\varepsilon^{2}}\log\tfrac{1}{\gamma}+n\log q\bigr{)}$.
We again emphasize that the continuous-setting algorithm can be used for integer distributions, and is often more efficient depending on the parameters. The problem ${P^{{\varepsilon}}_{\tt ratio-all}}$ was first considered in [16], under the name omnithermal approximation, along with an algorithm of cost $O(\frac{q^{2}}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$. A number of algorithms have been developed for pointwise estimation of $Q(\beta_{\max})$, which we denote by ${P^{{\varepsilon}}_{\tt ratio-pt}}$, with steadily improving sample complexities [4, 25, 14]; the best prior algorithm [14] had cost $O((q\log n)(\log q+\log\log n+\varepsilon^{-2}))$. No specialized algorithms were known for the integer setting.
We also show lower bounds for ${P^{{\varepsilon}}_{\tt ratio-pt}}$ (and, a fortiori, for ${P^{{\varepsilon}}_{\tt ratio-all}}$) for black-box sampling algorithms, which match Theorem 3 up to $\operatorname*{polylog}(n,q)$ factors:
Theorem 4.
Let $n\geq n_{0},q\geq q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0},\gamma<1/4$ for absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu$ which satisfy the given bounds $n$ and $q$ such that:
  • $\mu$ is continuous and ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega\bigl{(}\frac{q}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}$.
  • $\mu$ is log-concave and ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega\bigl{(}\frac{\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma}\bigr{)}.$
Explaining the problem definitions.
The definitions of the two problems ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$, and especially the precise form of the error term for ${P^{\delta,\varepsilon}_{\tt count}}$, may seem somewhat arbitrary. To illustrate, we note a few reductions showing how the error bounds for these problems can be combined to give estimates of the entire family of Gibbs distribution:
Theorem 5.
Given a solution for ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$, we can produce estimates
$\hat{\mu}_{\alpha}(x)=\frac{\hat{\pi}(x)e^{\alpha x}}{\hat{Q}(\alpha\mid\mathcal{D})}$
for any $\alpha\in[\beta_{\min},\beta_{\max}],x\in\mathcal{F}$ such that $|\hat{\mu}_{\alpha}(x)-\mu_{\alpha}(x)|\leq 3\varepsilon(\mu_{\alpha}(x)+\delta)$.
Theorem 6.
Given a solution for ${P^{1/n,\varepsilon/3}_{\tt count}}$ in the integer setting, we can solve ${P^{{\varepsilon}}_{\tt ratio-all}}$ with data structure $\mathcal{D}=\hat{\pi}$ and setting
$\hat{Q}(\alpha\mid\mathcal{D})=\sum_{x\in\mathcal{H}}\hat{\pi}(x)e^{\alpha x}.$
Indeed, our algorithms for ${P^{\delta,\varepsilon}_{\tt count}}$ in the general integer setting and log-concave setting immediately give us the corresponding algorithms in Theorem 3, essentially for free. See Appendix A for proofs.
1.2 Algorithmic sampling-to-counting
As a concrete application of our algorithms, consider the following scenario: we have a combinatorial system with weighted items, and we have an algorithm to (approximately) sample from the corresponding Gibbs distribution. The runtime (e.g. mixing time of a Markov chain) may depend on $\beta$. As a few prominent examples:
  1. 1.
    Connected subgraphs of a given graph; the weight of a subgraph is the number of edges [12, 11, 7].
  2. 2.
    Matchings of a given graph; the weight of a matching, again, is the number of edges [19, 17].
  3. 3.
    Independent sets in bounded-degree graphs; the weight is the size of the independent set [9, 17].
  4. 4.
    Vertex cuts in the ferromagnetic Ising model; the weight is the number of crossing edges [6].
A closely related statistical problem is to count the objects of a given weight class. There have been a number of problem-specific methods to use the sampling algorithm for this task, e.g. [18, 9]. Our algorithms for ${P_{\tt count}}$ can be combined with any prior sampling algorithms to yield improved and unified algorithmic results, essentially for free. As some examples, we will show the following:
Theorem 7.
Let $G=(V,E)$ be a connected graph, and for $i=0,\dots,|E|-|V|+1$ let $c_{i}$ be the number of connected subgraphs with $|E|-i$ edges. There is a fully-polynomial randomized approximation scheme (FPRAS) to simultaneously estimate all values $c_{i}$ to relative error $\varepsilon$ in time $\tilde{O}(|E|^{3}/\varepsilon^{2})$. (Throughout, we use the notation $\tilde{O}(x)=x\operatorname*{polylog}(x)$.)
Theorem 8.
Let $G=(V,E)$ be a graph and for each $i=0,\dots,|V|/2$ let $c_{i}$ be the number of matchings in $G$ with $i$ edges. Suppose $c_{|V|/2}>0$ and $c_{|V|/2-1}/c_{|V|/2}\leq f$ for a known parameter $f$. There is an FPRAS to simultaneously estimate all counts $c_{i}$ to relative error $\varepsilon$ in time $\tilde{O}(|E||V|^{3}f/\varepsilon^{2})$. In particular, if $G$ has minimum degree at least $|V|/2$, the time complexity is $\tilde{O}(|V|^{7}/\varepsilon^{2})$.
Theorem 9.
Let $D,\xi>0$ be arbitrary constants and let $G=(V,E)$ be a graph of maximum degree $D$. For $i=0,\dots,|V|$ let $c_{i}$ be the number of independent sets of size $i$. There is an FPRAS to simultaneously estimate all values $c_{0},\dots,c_{t}$ to relative error $\varepsilon$ in time $\tilde{O}(|V|^{2}/\varepsilon^{2})$ , where $t=(\alpha_{c}-\xi)|V|$ and $\alpha_{c}=\alpha_{c}(D)$ is the computational hardness threshold shown in [9].
Theorem 7 improves by a factor of $|E|$ over the algorithm in [12]. Similarly, Theorem 8 improves by a factor of $|V|$ over the algorithm in [19]. Theorem 9 matches the runtime of an FPRAS for a single count $c_{i}$ given in [17]. Full details of these constructions appear in Section 7.
There are two minor technical issues to clarify here. First, to get a randomized estimation algorithm, we must bound the computational complexity of our procedures in addition to the number of oracle calls. In all the algorithms we develop, the computational complexity is a small logarithmic factor times the query complexity. The computational complexity of the oracle is itself typically much larger than this overhead. Thus, our sampling procedures translate directly into efficient randomized algorithms, whose runtime is the expected sample complexity multiplied by the oracle’s computational complexity. We will not comment on computational issues henceforth.
Second, we may only have access to an approximate oracle $\tilde{\mu}_{\beta}$ instead of the exact oracle $\mu_{\beta}$. For example, given access to an MCMC sampler with stationary distribution $\mu_{\beta}$ and mixing time $\tau$, we can estimate $\mu_{\beta}$ to total variation distance $\rho$ by running for $\Theta(\tau\log(1/\rho))$ iterations. By a standard coupling argument, our results remain valid if exact oracles are replaced with sufficiently close approximate oracles (see e.g. [25, Remark 5.9]); see Appendix B for a formal statement.
1.3 Overview
We will develop two, quite distinct, types of algorithms: the first uses “cooling schedules” similar to [14], and the second is based on a new type of “covering schedule” for the integer setting. In Section 8, we show the lower bounds for the problems ${P_{\tt ratio-pt}}$ and ${P_{\tt count}}$.
Before the technical details, let us provide a high-level roadmap. For simplicity, we assume that tasks need to be solved with constant success probability.
The continuous setting.
For the problem ${P^{{\varepsilon}}_{\tt ratio-all}}$, we extend a statistical method of [14] called the Paired Product Estimator (see Section 3). This constructs a cooling schedule $\beta_{0},\beta_{1},\ldots,\beta_{t-1},\beta_{t}$, where the log partition function is near-linear between successive values $\beta_{i}$. As shown in [14], this property allows us to successively estimate the ratios $Q(\beta_{i})/Q(\beta_{i-1})$ via an unbiased estimator with low variance. In addition, we show that this property allows us to fill in values $Q(\alpha)$ for $\alpha\in(\beta_{i},\beta_{i-1})$ via linear interpolation.
For the problem ${P^{\delta,\varepsilon}_{\tt count}}$, we start with the following identity:
$\pi(y)=e^{-\beta y}\cdot\mu_{\beta}(y)\cdot Q(\beta)\qquad\text{for all $y\in\mathcal{F},\beta\in[\beta_{\min},\beta_{\max}]$}.$ (1)
We can estimate $Q(\beta)$ using our algorithm for ${P_{\tt ratio-all}}$, and we can estimate $\mu_{\beta}(y)$ by sampling. We then make use of the following important result: if $\mu_{\beta}([0,y])$ and $\mu_{\beta}([y,n])$ are both bounded below by constants, then $\mu_{\beta}(y)\geq\Omega(\Delta(y))$. For this value $\beta$, we can estimate $\mu_{\beta}(y)$ with $O(\frac{1}{\delta\varepsilon^{2}})$ samples, and from standard concentration bounds, the resulting estimate $\hat{\pi}$ satisfies the conditions of ${P^{\delta,\varepsilon}_{\tt count}}$. For example, if $\Delta(y)\geq\delta$, then $\mu_{\beta}(y)$, and hence $\pi(y)$, is within relative error $O(\varepsilon)$.
To estimate the full vector of counts, we find cut-points $x_{1}\geq x_{2}\geq\dots\geq x_{T}$, where each interval $[x_{i+1},x_{i}]$ has a corresponding value $\beta_{i}$ with $\mu_{\beta_{i}}([x_{i},n])\geq\Omega(1)$ and $\mu_{\beta_{i}}([0,x_{i+1}])\geq\Omega(1)$. Then every $y\in[x_{i+1},x_{i}]$ has $\mu_{\beta_{i}}(y)\geq\Omega(\Delta(y))$. In this way, a single value $\beta_{i}$ provides estimates for all the counts $c_{y}$ in the interval. We show that only $T=O(\sqrt{q\log n})$ distinct intervals are needed, leading to a cost of $O(\frac{\sqrt{q\log n}}{\delta\varepsilon^{2}})$ plus the cost of solving ${P_{\tt ratio-all}}$. The formal analysis appears in Section 4.
The integer setting.
To solve ${P^{\delta,\varepsilon}_{\tt count}}$, we develop a new data structure we call a covering schedule. This consists of a sequence $\beta_{\min}=\beta_{0},\beta_{1},\ldots,\beta_{t}=\beta_{\max}$ and corresponding values $k_{1},\dots,k_{t}$ so that $\mu_{\beta_{i}}(k_{i})$ and $\mu_{\beta_{i}}(k_{i+1})$ are large for all $i$. (The definition is adjusted slightly for the endpoints $i=0$ and $i=t$). Define $w_{i}=\min\{\mu_{\beta_{i}}(k_{i}),\mu_{\beta_{i}}(k_{i+1})\}$ (“weight” of $i$). If we take $\Omega(\frac{1}{w_{i}\varepsilon^{2}})$ samples from $\mu_{\beta_{i}}$, we can accurately estimate the quantities $\mu_{\beta_{i}}(k_{i}),\mu_{\beta_{i}}(k_{i+1})$, in turn allowing us to estimate
$\frac{Q(\beta_{i})}{Q(\beta_{i-1})}=e^{(\beta_{i}-\beta_{i-1})k_{i}}\frac{\mu_{\beta_{i-1}}(k_{i})}{\mu_{\beta_{i}}(k_{i})}$
By telescoping products, this in turn allows us to estimate every value $Q(\beta_{i})$.
Next, for each index $x\in\mathcal{H}$, we use binary search to find $\alpha$ with $\mu_{\alpha}([0,x])\approx\mu_{\alpha}([x,n])$ and then estimate $\mu_{\alpha}(x),\mu_{\alpha}(k_{i})$ by taking $O(\frac{1}{\delta\varepsilon^{2}})$ samples. If $\alpha$ lies in interval $[\beta_{i},\beta_{i+1}]$ of the covering schedule, we can use the estimates for $Q(\beta_{i})$ and $Q(\beta_{i+1})$ to estimate $Q(\alpha)$. (A slightly different procedure is used to handle exceptional cases where $\mu_{\alpha}(k_{i})$ is small.) Since we do this for each $j\in\mathcal{H}$, the overall cost of this second phase is roughly $O(\frac{n}{\delta\varepsilon^{2}})$.
In the log-concave setting, there is a more efficient algorithm for ${P^{\delta,\varepsilon}_{\tt count}}$. Here, for a fixed index $i$ of the covering schedule and for $j\in[k_{i},k_{i+1}]$ we have $\mu_{\beta_{i}}(j)\geq\min\{\mu_{\beta_{i}}(k_{i}),\mu_{\beta_{i}}(k_{i+1})\}=w_{i}$; thus, $\beta_{i}$ “covers” the interval $[k_{i},k_{i+1}]$. We can solve ${P^{\delta,\varepsilon}_{\tt count}}$ with $O(\frac{1}{\delta\varepsilon^{2}}+\sum_{i}\frac{1}{w_{i}\varepsilon^{2}})$ samples, by drawing $\Theta(\frac{1}{w_{i}\varepsilon^{2}})$ samples at $\beta_{i}$ and $\Theta(\frac{1}{\delta\varepsilon^{2}})$ samples at $\beta_{\min}$ and $\beta_{\max}$.
In either case, after solving ${P^{\delta,\varepsilon}_{\tt count}}$, we use Theorem 6 to solve ${P^{{\varepsilon}}_{\tt ratio-all}}$ for free.
Obtaining a covering schedule.
The general ${P_{\tt count}}$ algorithm above uses $O(\frac{n}{\delta\varepsilon^{2}}+\sum_{i}\frac{1}{w_{i}\varepsilon^{2}})$ samples, and similarly the log-concave algorithm uses $O(\frac{1}{\delta\varepsilon^{2}}+\sum_{i}\frac{1}{w_{i}\varepsilon^{2}})$ samples. We thus refer to the quantity $\sum_{i}\frac{1}{w_{i}}$ as the inverse weight of the schedule. In the most technically involved part of the paper, we produce a covering schedule with inverse weight $O(n\log n)$ (or $O(n)$ in the log-concave setting). Here we just sketch some key ideas.
First, we construct a “preschedule” where each interval can choose two different indices $\sigma^{-}_{i},\sigma^{+}_{i}$ instead of a single index $k_{i}$, with the indices interleaving as $\sigma_{i}^{-}\leq\sigma_{i+1}^{-}\leq\sigma_{i}^{+}\leq\sigma_{i+1}^{+}$. The algorithm repeatedly fill gaps: if some half-integer $\ell+1/2$ is not currently covered, we can select a value $\beta$ with $\mu_{\beta}([0,\ell])\approx\mu_{\beta}([\ell+1,n])$, along with corresponding values $\sigma^{-}\in[0,\ell],\sigma^{+}\in[\ell+1,n]$ with $\mu_{\beta}(\sigma^{+})\cdot(\sigma^{+}-\ell)\geq\Omega(\frac{1}{\log n})$ and $\mu_{\beta}(\sigma^{-})\cdot(\ell-\sigma^{-}+1)\geq\Omega(\frac{1}{\log n})$. The interval $[\sigma^{-},\sigma^{+}]$ then fills the gap and has weight $w\geq\Omega(\frac{1}{(\sigma^{+}-\sigma^{-})\log n})$.
At the end of the process, we throw away redundant intervals so that each $x$ is covered by at most two intervals, and “uncross” them into a schedule with $k_{i}\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$. Since $\frac{1}{w_{i}}\leq O((\sigma^{+}_{i}-\sigma^{-}_{i})\log n)$ for each $i$, this gives an $O(n\log n)$ bound of the inverse weight of the schedule.
2 Preliminaries
Define $z(\beta)=\log Z(\beta)$ and $z(\beta_{1},\beta_{2})=\log\frac{Z(\beta_{2})}{Z(\beta_{1})}=\log\frac{Q(\beta_{2})}{Q(\beta_{1})}$; note that $z(\beta_{\min},\beta_{\max})\leq q$ by definition. We write $z^{\prime}$ for the derivative of function $z$; observe (see [14]) that ${\mathbb{E}}_{X\sim\mu_{\beta}}[X]=z^{\prime}(\beta)$. As a consequence, $z^{\prime}$ is nonnegative everywhere and so the function $z$ is concave-up.
Define the Chernoff separation functions $F_{+}(u,t)=\bigl{(}\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\bigr{)}^{u}$ and $F_{-}(u,t)=\bigl{(}\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\bigr{)}^{u}$, where $\delta=t/u$. These are upper bounds on the probability that a binomial random variable with mean $u$ is larger than $u+t$ or smaller than $u-t$, respectively. We also write $F(u,t)=F_{+}(u,t)+F_{-}(u,t)$. We use some well-known bounds on these functions, e.g. $F_{+}(u,t)\leq e^{-t^{2}/(3u)}$ for $t\leq u$ and $F_{-}(u,t)\leq e^{-t^{2}/(2u)}$ for all $t$ and $F(u,t)\leq 2(1+t/u)^{-t/2}$ for all $t$.
For a random variable $X$, we write ${{\mathbb{S}}[{X}]}=\frac{{\mathbb{E}}[X^{2}]}{({\mathbb{E}}[X])^{2}}-1=\frac{\text{Variance}(X)}{({\mathbb{E}}[X])^{2}}$ for the relative variance of $X$.
We write $\mu_{\beta}(x,y),\mu_{\beta}[x,y)$ instead of $\mu_{\beta}((x,y)),\mu_{\beta}([x,y))$, etc. for readability. For any fixed value $\chi$, the map $\beta\mapsto\mu_{\beta}[0,\chi)$ is a non-increasing function of $\chi$, since $\frac{\mu_{\beta}[\chi,n]}{1-\mu_{\beta}[\chi,n]}=\frac{\sum_{x\geq\chi}c_{x}e^{\beta x}}{\sum_{x<\chi}c_{x}e^{\beta x}}$; when $\beta$ increases by $\delta$, the numerator is multiplied by at least $e^{\delta\chi}$ and the denominator is multiplied by at most $e^{\delta\chi}$. At several places, we also make use of the following inequality:
for values $\text{for values $\alpha\leq\beta,x\leq y$:}\qquad{\mu_{\alpha}(x)\mu_{\beta}(y)}\geq\mu_{\alpha}(y)\mu_{\beta}(x)$ (Uncrossing Inequality)
2.1 The Balance subroutine
Given a target $\chi$, we need to find a value $\beta$ with $\mu_{\beta}[0,\chi]\approx 1/2\approx\mu_{\beta}[\chi,n]$. That is, $\chi$ is approximately the median of the distribution $\mu_{\beta}$. We can find the target value $\beta$ via noisy binary search.
Formally, for $\tau\in(0,1/2),\chi\in[0,n]$, we define ${\tt BalancingVals}(\chi,\tau)$ to be the set of values $\beta\in[\beta_{\min},\beta_{\max}]$ satisfying the following two properties:
  • Either $\beta=\beta_{\min}$ or $\mu_{\beta}[0,\chi]\geq\tau$
  • Either $\beta=\beta_{\max}$ or $\mu_{\beta}[\chi,n]\geq\tau$
The following observation explains the motivation for the definition.
Proposition 10.
If $\beta\in{\tt BalancingVals}(x,\tau)$, then $\mu_{\beta}(x)\geq\tau\Delta(x)$.
Proof.
Consider $\alpha\in[\beta_{\min},\beta_{\max}]$ with $\mu_{\alpha}(x)=\Delta(x)$. The result is clear if $\alpha=\beta$. Suppose that $\alpha<\beta$; the case $\alpha>\beta$ is completely analogous. So $\beta>\beta_{\min}$, and since $\beta\in{\tt BalancingVals}(x,\tau)$, this implies $\mu_{\beta}[0,x]\geq\tau$. We then have:
$\mu_{\alpha}(x)=\frac{c_{x}e^{\alpha x}}{\sum_{y}c_{y}e^{\alpha y}}\leq\frac{c_{x}e^{\alpha x}}{\sum_{y\leq x}c_{y}e^{\alpha y}}\leq\frac{c_{x}e^{\beta x}}{\sum_{y\leq x}c_{y}e^{\beta y}}=\frac{\mu_{\beta}(x)}{\mu_{\beta}[0,x]}\leq\frac{\mu_{\beta}(x)}{\tau}.\qed$
In Appendix D, we show the following result:
Theorem 11.
Let $\tau\in(0,1/2)$ be an arbitrary constant. Then subroutine ${\tt Balance}(\chi,\tau,\gamma)$ has cost $O(\log\frac{nq}{\gamma})$. With probability at least $1-\gamma$, it returns a value $\beta\in{\tt BalancingVals}(\chi,\tau)$ (we say in this case that the call is good).
The algorithm can be accelerated if we can solve ${P_{\tt ratio-all}}$ as a preprocessing step:
Theorem 12.
Let $\tau\in(0,1/2)$ be an arbitrary constant and let $\varepsilon=(1/2-\tau)/30$. If we are given a data structure $\mathcal{D}$ for ${P^{{\varepsilon}}_{\tt ratio-all}}$, then subroutine ${\tt BalancePreprocessed}(\chi,\tau,\mathcal{D},\gamma)$ has cost $O(\log\frac{q}{\gamma})$. If $\mathcal{D}$ is $\varepsilon$-close, then with probability at least $1-\gamma$, it returns a value $\beta\in{\tt BalancingVals}(\chi,\tau)$.
2.2 Statistical sampling
We discuss a few procedures for statistical sampling in our algorithms; see Appendix C for proofs.
First, we can obtain an unbiased estimator of the probability vector $\mu_{\beta}$ by computing empirical frequencies $\hat{\mu}_{\beta}$ from $N$ independent samples from $\mu_{\beta}$; we denote this process as $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta,N)$. We record the following standard concentration bound, which we will use repeatedly:
Lemma 13.
[…]
In particular, if Eq. (2) holds and $\max\{p,\hat{p}\}\geq\bar{p}$, then $|\log\hat{p}-\log p|\leq\varepsilon$.
Many of our algorithms are based on calling $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta,N)$ and making decisions depending on the values $\hat{\mu}_{\beta}(I)$ for certain sets $I\subseteq\mathcal{F}$; they succeed when the estimates $\hat{\mu}_{\beta}(I)$ are close to $\mu_{\beta}(I)$. We say the execution of Sample well-estimates $I$ if Eq. (2) hold for $p=\mu_{\beta}(I)$ and $\hat{p}=\hat{\mu}_{\beta}(I)$; otherwise it mis-estimates $I$. Likewise we say Sample well-estimates $k$ if it well-estimates the singleton set $I=\{k\}$. Since this comes up so frequently, we write
$\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta;\varepsilon,\bar{p},\gamma)$
as shorthand for $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta,\big{\lceil}\frac{3e^{\varepsilon}\log(4/\gamma)}{(1-e^{-\varepsilon})^{2}\bar{p}}\big{\rceil})$. It has cost $O(\frac{\log(1/\gamma)}{\varepsilon^{2}\bar{p}})$, and each individual set $I$ is well-estimated with probability at least $1-\gamma$. Here $\varepsilon$ should be interepreted as the desired relative error, $\bar{p}$ as the minimum “detection threshold” for $\mu_{\beta}$, and $\gamma$ as the desired failure probability.
In a similar way, our algorithms for ${P^{\delta,\varepsilon}_{\tt count}}$ estimate $\pi(x)$ by sampling $\hat{\mu}_{\alpha}(x)$ for a well-chosen value $\alpha$. We record the following straightforward error bound:
Lemma 14.
Suppose that for $x\in\mathcal{F}$, we have $\alpha\in[\beta_{\min},\beta_{\max}]$ and non-negative parameters $\hat{Q}(\alpha),\hat{\mu}_{\alpha}(x)$ (which may depend upon $x$), satisfying the following bounds:
(A1) $|\log\hat{Q}(\alpha)-\log Q(\alpha)|\leq\varepsilon/3$.
(A2) $|\hat{\mu}_{\alpha}(x)-\mu_{\alpha}(x)|\leq\varepsilon/3\cdot\mu_{\alpha}(x)(1+\delta/\Delta(x))$.
Then the estimated value $\hat{\pi}(x)=\hat{Q}(\alpha)e^{-\alpha x}\hat{\mu}_{\alpha}(x)$ satisfies the criteria for the problem ${P^{\delta,\varepsilon}_{\tt count}}$.
Finally, in a few places, we need to estimate certain telescoping products. Direct Monte Carlo sampling does not give strong tail bounds, so we use a standard method based on median amplification.
Theorem 15.
Suppose we can draw independent samples from non-negative random variables $X_{1},\dots,X_{N}$. Given input $\gamma\in(0,1/2)$ and $\tau\geq 1$, the procedure ${\tt EstimateProducts}(X,\tau,\gamma)$ returns a vector of estimates $(\hat{X}^{\tt prod}_{1},\dots,\hat{X}^{\tt prod}_{N})$ using $O(N\tau\log\tfrac{1}{\gamma})$ total samples of the $X$ variables. If $\tau\geq\sum_{i=1}^{N}{{\mathbb{S}}[{X_{i}}]}$, then with probability at least $1-\gamma$, it holds that
for all $\text{for all $i=1,\dots,N$: }\frac{\hat{X}_{i}^{\tt prod}}{\prod_{j=1}^{i}{\mathbb{E}}[X_{j}]}\in[e^{-\varepsilon},e^{\varepsilon}]\qquad\qquad\text{for $\varepsilon=\sqrt{\sum_{i=1}^{N}{{\mathbb{S}}[{X_{i}}]}/\tau}\in(0,1)$}$
In this context, we also define $\hat{X}^{\tt prod}_{0}=1$.
3 Solving ${P^{{\varepsilon}}_{\tt ratio-all}}$ in the continuous setting
In this section, we will develop our algorithm PratioContinuous to solve ${P^{{\varepsilon}}_{\tt ratio-all}}$ based on a stochastic process called TPA developed in [14]. The TPA process generates a sequence $\beta_{i}$ of values in the range $[\beta_{\min},\beta_{\max}]$ as follows:
1 set $\beta_{0}=\beta_{\max}$
2 for $i=1$ to +$\infty$+ ∞ do
3       sample $K\sim\mu_{\beta_{i-1}}$ and sample $U\sim\text{Unif}([0,1])$.
4       set $\beta_{i}=\beta_{i-1}+\frac{\log U}{K}$
5       if $K=0$ or $\beta_{i}<\beta_{\min}$ then return set $B=\{\beta_{1},\dots,\beta_{i-1}\}$
6
Algorithm 1 The TPA process. Output: a multiset of values in the interval $[\beta_{\min},\beta_{\max}]$
${\tt TPA}(k)$ is defined to be the union of $k$ independent runs of Algorithm 1. The critical property shown in [15, 16] is that $z({\tt TPA}(k))$ is a rate-$k$ Poisson Point Process on $[z(\beta_{\min}),z(\beta_{\max})]$. In other words, if $\{\beta_{1},\dots,\beta_{\ell}\}$ is the output of ${\tt TPA}(k)$, then the random variables $z_{i}=z(\beta_{i})$ are generated by the following process:
1 set $z_{0}=z(\beta_{\max})$
2 for $i=1$ to +$\infty$+ ∞ do
3       draw $\eta$ from the Exponential distribution of rate $k$, and set $z_{i}=z_{i-1}-\eta$
4       if $z_{i}<z(\beta_{\min})$ then return set $\{z_{1},\dots,z_{i-1}\}$
5
Algorithm 2 Equivalent process for generating $z({\tt TPA}(k))$.
As noted in [16], the TPA process can be used for a simple algorithm for ${P^{{\varepsilon}}_{\tt ratio-all}}$, which already gives us half of our result.
compute $B\leftarrow{\tt TPA}(k)$ and set $\mathcal{D}=B$.     Query algorithm $Q(\alpha\mid{\cal D})$:
return $\hat{Q}(\alpha|{\cal D})=\exp\bigl{(}|B\cap[\beta_{\min},\alpha)|/k\bigr{)}$
Algorithm 3 Algorithm $\tt{TPA-Pratio}(k)$:
Theorem 16 ([16]).
Algorithm TPA-Pratio has cost $O(k(1+\log Q(\beta_{\max})))$. If $k\geq\frac{20\log Q(\beta_{\max})}{\varepsilon^{2}}\log\tfrac{3}{\gamma}$, then the data structure $\mathcal{D}$ is $\varepsilon$-close with probability at least $1-\gamma$.
In particular,with $k=\lceil\frac{20q}{\varepsilon^{2}}\log\tfrac{3}{\gamma}\rceil$, then the algorithm solves PratioAll with cost $O(\frac{q^{2}}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$.
Proof.
The runtime follows since each run of TPA has cost $O(1+\log Q(\beta_{\max}))$.
For any values $\beta_{1}<\beta_{2}$, let $B(\beta_{1},\beta_{2})=|B\cap(\beta_{1},\beta_{2})|$. This is a Poisson random variable with mean $k\cdot z(\beta_{1},\beta_{2})$. So ${{\mathbb{P}}\!\bm{\left[}|B(\beta_{1},\beta_{2})/k-z(\beta_{1},\beta_{2})|\geq\varepsilon/2\bm{}}\leq F(qk,\varepsilon k/2)\leq 2e^{-\frac{\varepsilon^{2}k}{12q}}\leq\gamma/2$ for fixed $\beta_{1},\beta_{2}$.
Let $\mathcal{E}$ denote the event that $\mathcal{D}$ fails to be good for some value $\alpha$. Consider the random process where $B\cap[\beta_{\min},\alpha]$ is revealed while $\alpha$ gradually increases; suppose $\mathcal{E}$ occurs and let $\alpha$ be the first value with $|B(\beta_{\min},\alpha)/k-z(\beta_{\min},\alpha)|>\varepsilon$. Since $z(B)$ is a Poisson process, $B\cap(\alpha,\beta_{\max})$ retains its original, unconditioned probability distribution. Accordingly, there is a probability of at least $21-\gamma/$ 2 that $|B(\alpha,\beta_{\max})/k-z(\alpha,\beta_{\max})|\leq\varepsilon/2$, in which case we would get
$|B(\beta_{\min},\beta_{\max})/k-z(\beta_{\min},\beta_{\max})|=|(B(\beta_{\min},\alpha)/k-z(\beta_{\min},\alpha))+(B(\alpha,\beta_{\max})/k-z(\alpha,\beta_{\max}))|$
$\qquad\geq|B(\beta_{\min},\alpha)/k-z(\beta_{\min},\alpha)|-|B(\alpha,\beta_{\max})/k-z(\alpha,\beta_{\max})|\geq\varepsilon-\varepsilon/2\geq\varepsilon/2.$
[…]
3.1 Paired Product Estimator
To improve the sample complexity to $O(\frac{q\log n}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$, we need a more advanced statistical technique called the Paired Product Estimator developed in [14]. Here, we describe a simple version of the algorithm. [^2] This is a simplified version of the PPE algorithm of [14]. See also [14, 22] for a different, stand-alone version of the PPE algorithm to directly solve ${P^{{\varepsilon}}_{\tt ratio-pt}}$ with high probability.
1 compute $B\leftarrow{\tt TPA}(k)$ sorted as $B=\{\beta_{1},\dots,\beta_{t-1}\}$
2 define $\beta_{0}=\beta_{\min}$ and $\beta_{t}=\beta_{\min}$.
3 for $i=1,\ldots,t$ do
4       define random variable $W_{i}=\exp\bigl{(}\frac{\beta_{i}-\beta_{i-1}}{2}\cdot K\bigr{)}$ where $K$ is drawn from $\mu_{\beta_{i-1}}$
5       define random variable $V_{i}=\exp\bigl{(}-\frac{\beta_{i}-\beta_{i-1}}{2}\cdot K\bigr{)}$ where $K$ is drawn from $\mu_{\beta_{i}}$
6set $\hat{W}^{\tt prod}={\tt EstimateProducts}(W_{i},32,1/10)$
7 set $\hat{V}^{\tt prod}={\tt EstimateProducts}(V_{i},32,1/10)$
return ${\cal D}=((\beta_{0},\ldots,\beta_{t}),(\hat{Q}(\beta_{0}),\ldots,\hat{Q}(\beta_{t}))$ where $\hat{Q}(\beta_{i})=\hat{W}^{\tt prod}_{i}/\hat{V}^{\tt prod}_{i}$     Query algorithm $Q(\alpha\mid{\cal D})$:
8 find index $i$ with $\alpha\in[\beta_{i},\beta_{i+1}]$
9 let $\alpha=(1-v)\beta_{i-1}+v\beta_{i}$ for $v\in[0,1]$
return $\hat{Q}(\alpha|{\cal D})=\exp\bigl{(}(1-v)\log\hat{Q}(\beta_{i-1})+v\log\hat{Q}(\beta_{i})\bigr{)}$
Algorithm 4 Algorithm ${\tt PPE-Pratio}(k)$
In this context, the sequence $(\beta_{0},\dots,\beta_{t})$ constructed at Line 2 is called a cooling schedule. We show the following main result for this algorithm:
Lemma 17.
The algorithm PPE-Pratio has cost $O(kq)$. If $k\geq\frac{10\log(z^{\prime}(\beta_{\max})/z^{\prime}(\beta_{\min}))}{\varepsilon^{2}}$ for $\varepsilon\in(0,1)$, then the resulting data structure $\mathcal{D}$ is $\varepsilon$-close with probability at least $0.7$.
For the complexity bound, Line 1 has cost $O(kq)$. The two applications of EstimateProducts have cost $O(t)$ given the cooling schedule $\beta_{i}$, and the expected size of $t$ is $O(kq)$.
For the correctness bound, we consider curvature parameters $\kappa_{i}$ and $\kappa$ of the schedule defined as:
$\kappa_{i}=z(\beta_{i-1})-2z\bigl{(}\tfrac{\beta_{i-1}+\beta_{i}}{2}\bigr{)}+z(\beta_{i}),\qquad\qquad\kappa=\sum_{i=1}^{t}\kappa_{i}$
Our main technical result here is the following bound:
Lemma 18.
There holds ${\mathbb{E}}[\kappa]\leq\frac{\log(z^{\prime}(\beta_{\max}))-\log(z^{\prime}(\beta_{\min}))}{k}.$
Proof.
As shown in [14, Lemma 3.2], we have $\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\geq\exp(\frac{2\kappa_{i}}{z(\beta_{i-1},\beta_{i})})$ for each $i$ (see also [25]). (Here $z^{\prime}(\beta)$ denotes the derivative of function $z$ at value $\beta$). Equivalently, we have $\kappa_{i}\leq\frac{z(\beta_{i-1},\beta_{i})}{2}\cdot\log\bigl{(}\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\bigr{)}$.
Observe that $z(\beta)$ and $z^{\prime}(\beta)$ are both strictly increasing functions of $\beta$. Thus, for any fixed value $x\in[z^{\prime}(\beta_{\min}),z^{\prime}(\beta_{\max})]$, there is a unique value $\beta\in[\beta_{\min},\beta_{\max}]$ with $z^{\prime}(\beta)=x$, and a unique index $i$ with $z(\beta)\in[z(\beta_{i-1}),z(\beta_{i}))$. Define $A_{x}=z(\beta_{i-1},\beta_{i})$, that is, the distance between the nearest points of the cooling schedule. We can rewrite the bound for $\kappa_{i}$ then as
$\kappa_{i}\leq\frac{z(\beta_{i-1},\beta_{i})}{2}\cdot\log\bigl{(}\frac{z^{\prime}(\beta_{i})}{z^{\prime}(\beta_{i-1})}\bigr{)}=\frac{1}{2}\int_{z^{\prime}(\beta_{i-1})}^{z^{\prime}(\beta_{i})}\frac{A_{x}}{x}\ \operatorname{d}\!{x}$
so that:
$\kappa\leq\frac{1}{2}\sum_{i}\int_{z^{\prime}(\beta_{i-1})}^{z^{\prime}(\beta_{i})}\frac{A_{x}}{x}\ \operatorname{d}\!{x}=\frac{1}{2}\int_{z^{\prime}(\beta_{\min})}^{z^{\prime}(\beta_{\max})}\frac{A_{x}}{x}\ \operatorname{d}\!{x}.$ (3)
We claim that, for any fixed $x$, we have ${\mathbb{E}}[A_{x}]\leq 2/k$. For, suppose $z^{\prime}(\beta^{\ast})=x$ and let $z^{\ast}=z(\beta^{\ast})$. Consider a Poisson process $X_{1},X_{2},\ldots$ on $[z^{\ast},+\infty)$ and a Poisson process $X_{-1},X_{-2},\ldots$ on $(-\infty,z^{\ast}]$, both with rate $k$. By the superposition theorem for Poisson processes [21, page 16], the bidirectional sequence ${\bf X}=\ldots,X_{-2},X_{-1},X_{1},X_{2},\ldots$ is also a rate-$k$ Poisson process on (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). As discussed earlier, the sequence $z(\beta_{i})$ has the same distribution as the subsequence ${\bf X}\cap[z(\beta_{\min}),z(\beta_{\max})]$. So
${\mathbb{E}}[A_{x}]={\mathbb{E}}[\min\{z(\beta_{\max}),X_{1}\}-\max\{z(\beta_{\min}),X_{-1}\}]\leq{\mathbb{E}}[X_{1}-z^{\ast}]+{\mathbb{E}}[z^{\ast}-X_{-1}]=2/k$
where the last equality uses the fact that $X_{1}-z^{\ast}$ and $z^{\ast}-X_{-1}$ are rate-$k$ Exponential random variables.
Substituting into Eq. (3) gives:
${\mathbb{E}}[\kappa]\leq\frac{1}{2}\int_{z^{\prime}(\beta_{\min})}^{z^{\prime}(\beta_{\max})}\frac{{\mathbb{E}}[A_{x}]}{x}\ \operatorname{d}\!{x}\leq\frac{1}{2}\int_{z^{\prime}(\beta_{\min})}^{z^{\prime}(\beta_{\max})}\frac{2}{kx}\ \operatorname{d}\!{x}=\frac{\log(z^{\prime}(\beta_{\max}))-\log(z^{\prime}(\beta_{\min}))}{k}.\qed$
Proposition 19.
If $k\geq\frac{10(\log z^{\prime}(\beta_{\max})-\log z^{\prime}(\beta_{\min}))}{\varepsilon^{2}}$ for $\varepsilon\in(0,1)$, then with probability at least $0.7$, the following two bounds hold: (a) $\kappa\leq\varepsilon^{2}$ and (b) $|\log(\hat{W}_{i}^{\tt prod}/\hat{V}_{i}^{\tt prod})-\log Q(\beta_{i})|\leq\varepsilon/2$ for all $i=0,\dots,t$.
Proof.
By Markov’s inequality, we have $\kappa\leq\varepsilon^{2}$ with probability at least $0.9$. Suppose this holds, and condition on the fixed schedule $B$. Denote $\bar{\beta}_{i-1,i}=\frac{\beta_{i-1}+\beta_{i}}{2}$ for each $i$. A calculation shows (see [14]):
${\mathbb{E}}[W_{i}]=\frac{Z(\bar{\beta}_{i-1,i})}{Z(\beta_{i-1})}\;\;\quad\quad{\mathbb{E}}[V_{i}]=\frac{Z(\bar{\beta}_{i-1,i})}{Z(\beta_{i})}\;\;\quad$ ${{\mathbb{S}}[{W_{i}}]}={{\mathbb{S}}[{V_{i}}]}=\frac{Z(\beta_{i-1})Z(\beta_{i})}{Z(\bar{\beta}_{i-1,i})^{2}}-1=e^{\kappa_{i}}-1$
So $\sum_{i}{{\mathbb{S}}[{W_{i}}]}\leq\sum_{i}(e^{\kappa_{i}}-1)\leq e^{\sum_{i}\kappa_{i}}-1=e^{\kappa}-1$; since $\kappa\leq\varepsilon^{2}\leq 1$, this is at most $2\varepsilon^{2}$. Likewise $\sum_{i}{{\mathbb{S}}[{V_{i}}]}\leq 2\varepsilon^{2}$. So by Theorem 15, with probability at least $0.8$ the values $\hat{W}^{\tt prod},\hat{V}^{\tt prod}$ returned by EstimateProducts satisfy
$\frac{\hat{V}^{\tt prod}_{i}}{\prod_{j=1}^{i}{\mathbb{E}}[V_{j}]}\in[e^{-\varepsilon^{\prime},\varepsilon^{\prime}}],\qquad\frac{\hat{W}^{\tt prod}_{i}}{\prod_{j=1}^{i}{\mathbb{E}}[W_{j}]}\in[e^{-\varepsilon^{\prime},\varepsilon^{\prime}}]\qquad\text{for all $i=0,\dots,t$}$
where $\varepsilon^{\prime}=\varepsilon^{2}/4$. In this case, each value $\hat{Q}(\beta_{i})=\hat{W}^{\tt prod}_{i}/\hat{V}^{\tt prod}_{i}$ is within $[e^{-\varepsilon/2},e^{\varepsilon/2}]$ of the product $\prod_{j=1}^{i}\frac{{\mathbb{E}}[W_{j}]}{{\mathbb{E}}[V_{j}]}$; by telescoping products, this is precisely $Z(\beta_{i})/Z(\beta_{0})=Q(\beta_{i})$. ∎
Proposition 20.
If the bounds of Proposition 19 hold, then the data structure $\mathcal{D}$ is $\varepsilon$-close.
Proof.
Consider $\alpha=(1-v)\beta_{i-1}+v\beta_{i}$ for $v\in[0,1]$. Suppose that $v\leq 1/2$; the case where $v\geq 1/2$ is completely symmetric. We need to show:
$|(1-v)\hat{z}(\beta_{i-1})+v\hat{z}(\beta_{i})-z(\alpha)|\leq\varepsilon$ (4)
where we write $\hat{z}(\beta)=\log\hat{Q}(\alpha\mid\mathcal{D})$ for brevity. We can rewrite the definition of $\kappa_{i}$ as:
$\kappa_{i}=2z(\bar{\beta},\beta_{i})-z(\beta_{i-1},\beta_{i})\qquad\qquad\text{for $\bar{\beta}=\tfrac{\beta_{i-1}+\beta_{i}}{2}$}$ (5)
Since $z$ is an increasing concave-up function, $z(\beta_{i-1},\alpha)\leq v\cdot z(\beta_{i-1},\beta_{i})$ and $z(\beta_{i-1},\alpha)\geq z(\beta_{i-1},\beta_{i})-2(1-v)z(\bar{\beta},\beta_{i})$. By Eq. (5), this implies that $z(\beta_{i-1},\alpha)\geq v\cdot z(\beta_{i-1},\beta_{i})-\kappa_{i}$. So overall
$|z(\beta_{i-1},\alpha)-v\cdot z(\beta_{i-1},\beta_{i})|=|(1-v)z(\beta_{i-1})+v\cdot z(\beta_{i})-z(\alpha)|\leq\kappa_{i};$
by Proposition 19(a) this at most $\varepsilon^{2}\leq\varepsilon/2$. By Proposition 19(b), we have $|\hat{z}(\beta_{i-1})-z(\beta_{i-1})|<\varepsilon/2$ and $|\hat{z}(\beta_{i})-z(\beta_{i})|<\varepsilon/2$, so this implies Eq. (4). ∎
This concludes the proof for Lemma 17.
3.2 A hybrid algorithm
Our hybrid algorithm for ${P^{{\varepsilon}}_{\tt ratio-all}}$ combines the two simpler algorithms in their appropriate parameter regimes.
1 if $q>\log n$ then
2       set $\beta_{\circ}\leftarrow\beta_{\max}$ and $t=0$
3       call $\mathcal{D}_{0}={\tt TPA-Pratio}(k_{1})$ for $k_{1}=\lceil\frac{20q}{\varepsilon^{2}}\log\tfrac{3}{\gamma}\rceil$
4else if $q\leq\log n$ then
5       set $\beta_{\circ}\leftarrow{\tt Balance}(1/2,1/4,\gamma/4)$
6       call $\mathcal{D}_{0}={\tt TPA-Pratio}(k_{1})$ with $k_{1}=\lceil\frac{40\cdot\log 4}{\varepsilon^{2}}\log\tfrac{12}{\gamma}\rceil$ for the interval $[\beta_{\min},\beta_{\circ}]$.
7       for $i=1,\dots,t=40\log\tfrac{2}{\gamma}$ do
8             Set $\mathcal{D}_{i}\leftarrow{\tt PPE-Pratio}(k_{2},\varepsilon/2)$ with $k_{2}=\lceil\frac{10\log(4n)}{\varepsilon^{2}}\rceil$ for the interval $[\beta_{\circ},\beta_{\max}]$
9
10
output tuple $\mathcal{D}=(\beta_{\circ},\mathcal{D}_{1},\dots,\mathcal{D}_{t})$     Query algorithm $Q(\alpha\mid{\cal D})$:
11 if $\alpha\leq\beta_{\circ}$ then
12       return $\hat{Q}(\alpha\mid\mathcal{D}_{0})$.
13else
14       return $\hat{Q}(\beta_{\circ}\mid\mathcal{D}_{0})\cdot\text{median}\{\hat{Q}(\alpha\mid\mathcal{D}_{1}),\dots,\hat{Q}(\alpha\mid\mathcal{D}_{t})\}$
Algorithm 5 Algorithm ${\tt PratioContinuous}(\varepsilon,\gamma)$
Theorem 21.
Algorithm PratioContinuous solves ${P^{{\varepsilon}}_{\tt ratio-all}}$ with cost
$O\Bigl{(}\frac{\min\{q^{2},q\log n\}}{\varepsilon^{2}}\log\tfrac{1}{\gamma}\Bigr{)}.$
Proof.
If $q>\log n$, then this simply reduces to calling TPA-Pratio and the result follows from Theorem 16. So suppose $q\leq\log n$. In this case, Line 5 has cost $O(\log\frac{nq}{\gamma})$ and each iteration of the loop at Line 7 has cost $O(k_{2}q)=O(\frac{q\log n}{\varepsilon^{2}})$. Line 6 has cost $O(k_{1}(1+\log Q(\beta_{\circ})))\leq O(\frac{q}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$. Overall, in this case the algorithm has cost $O(\frac{q\log n}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$.
For the algorithm correctness, suppose the call to Balance at Line 5 is good, which occurs with probability at least $41-\gamma/$ 4. In this case, we claim that the following two bounds hold:
$Q(\beta_{\circ})\leq 4,\qquad\qquad z^{\prime}(\beta_{\max})/z^{\prime}(\beta_{\circ})\leq 4n$ (6)
The first bound is immediate if $\beta_{\circ}=\beta_{\min}$. Otherwise, we have $\mu_{\beta_{\circ}}(0)=\mu_{\beta_{\circ}}[0,1/2]\geq 1/4$. We can estimate $Z(\beta_{\min})\geq c_{0}$ and $Z(\beta_{\circ})=\frac{c_{0}}{\mu_{\beta_{\circ}}(0)}\leq 4c_{0}$. So $Q(\beta_{\circ})\leq 4$.
Likewise, the second bound is immediate if $\beta_{\circ}=\beta_{\max}$. Otherwise, we have $\mu_{\beta_{\circ}}[1,n]\geq 1/4$. Thus $z^{\prime}(\beta_{\circ})={\mathbb{E}}_{X\sim\mu_{\beta_{\circ}}}[X]\geq\mu_{\beta_{\circ}}[1,n]\geq 1/4$. So $z^{\prime}(\beta_{\max})/z^{\prime}(\beta_{\circ})\leq n/(1/4)=4n$.
By Eq. (6), Theorem 16 and our choice of $k_{1}\geq\frac{80\log 4}{(\varepsilon/2)^{2}}\log\tfrac{3}{\gamma/4}$, with probability at least $41-\gamma/$ 4, data structure $\mathcal{D}_{0}$ is $\varepsilon/2$-close for the range $[\beta_{\min},\beta_{\circ}]$. In particular, $\hat{Q}(\beta_{\circ}\mid\mathcal{D}_{0})$ is within $e^{\pm\varepsilon/2}$ of $Q(\beta_{\circ})$. Also, by Lemma 17, each $\mathcal{D}_{i}$ is $\varepsilon/2$-close for the range $[\beta_{\circ},\beta_{\max}]$ with probability at least $0.7$. So with probability at least $21-F_{-}(0.7t,0.2t)\geq 1-\gamma/$ 2, at least half the structures $\mathcal{D}_{i}$ are $\varepsilon/2$-close; when this holds, then $\text{median}(\hat{Q}(\alpha\mid\mathcal{D}_{1}),\dots,\hat{Q}(\alpha\mid\mathcal{D}_{t}))$ is within $e^{\pm\varepsilon/2}$ of $\frac{Q(\alpha)}{Q(\beta_{\circ})}$ for all values $\alpha\in[\beta_{\circ},\beta_{\max}]$. Then the final data structure $\mathcal{D}$ is $\varepsilon$-close for the entire range $[\beta_{\min},\beta_{\max}]$. ∎
We have established the continuous-setting algorithm of Theorem 3.
4 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the continuous setting
In this section, we develop our algorithm for ${P^{\delta,\varepsilon}_{\tt count}}$, using the algorithm for ${P_{\tt ratio-all}}$ as a subroutine.
1 set $\mathcal{D}\leftarrow{\tt PratioContinuous}(\min\{1/120,\varepsilon/3\},\gamma/4)$.
2 initialize $x_{0}\leftarrow n$
3 for $t=1$ to $T=2\lceil\min\{q/2,\sqrt{q\log n}\}\rceil+4$  do
4       set $\beta_{t}\leftarrow{\tt BalancePreprocessed}(x_{t-1},1/4,\mathcal{D},\frac{\gamma}{10T})$
5       set $\hat{\mu}_{\beta_{t}}\leftarrow{\tt Sample}(\beta_{t},\lceil\frac{10^{6}\log{\frac{10T}{\delta\gamma}}}{\delta\varepsilon^{2}}\rceil)$
6       if $\beta_{t}>\beta_{\min}$ then
7             set $x_{t}$ to be the minimum value with $\hat{\mu}_{\beta_{t}}[0,x_{t}]\geq 1/90$
8             foreach $y\in(x_{t},x_{t-1}]$ do set $\hat{\pi}(y)=\hat{Q}(\beta_{t}\mid\mathcal{D})e^{-\beta_{t}}\hat{\mu}_{\beta}(y)$
9      else if $\beta_{t}=\beta_{\min}$ then
10             foreach $y\in[0,x_{t-1}]$ do set $\hat{\pi}(y)=e^{-\beta_{\min}}\hat{\mu}_{\beta_{\min}}(y)$
11             return all estimated values $\hat{\pi}$
12
output ERROR and terminate
Algorithm 6 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ for error parameter $\gamma$.
Theorem 22.
Algorithm 6 solves ${P^{\delta,\varepsilon}_{\tt count}}$ with cost
$O\Bigl{(}\min\{\frac{q^{2}\log\tfrac{1}{\gamma}+(q/\delta)\log\tfrac{q}{\delta\gamma}}{\varepsilon^{2}},\frac{q\log n\log\tfrac{1}{\gamma}+(\sqrt{q\log n}/\delta)\log\tfrac{q}{\delta\gamma}}{\varepsilon^{2}}\}\Bigr{)}$
The complexity bound follows immediately from specification of subroutines. We next analyze the success probability. Throughout, we suppose that that the call at Line 1 succeeds, and produces a good data structure $\mathcal{D}$. There are a number of intermediate calculations.
Proposition 23.
With probability at least $21-\gamma/$ 2, the following conditions hold for all iterations $t$:
(a) If $\beta_{t}<\beta_{\max}$, then $\mu_{\beta_{t}}[x_{t-1},n]\geq 1/4$.
(b) If $\beta_{t}>\beta_{\min}$, then $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$ and $x_{t}<x_{t-1}$ and $\mu_{\beta_{t}}[0,x_{t})\leq 1/80$ and $\mu_{\beta_{t}}[0,x_{t}]\geq 1/100$.
Proof.
With probability at least $1-\frac{\gamma}{10T}$, the call to Balance at each iteration $t$ is good. Let us suppose this holds. By definition, this shows (a). For (b), suppose $\beta_{t}>\beta_{\min}$. By specification of Balance we immediately have $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$. Let $v,w$ be the minimum values with $\mu_{\beta_{t}}[0,v]\geq 1/80$ and $\mu_{\beta_{t}}[0,w]\geq 1/100$ respectively. (These both exist since the function $x\mapsto\mu_{\beta}[0,x]$ is right-continuous.) Then $\mu_{\beta_{t}}[0,v)\leq 1/80,\mu_{\beta_{t}}[0,w)\leq 1/100$. Since $\mu_{\beta_{t}}[0,x_{t-1})\geq 1/4$, we have $w\leq v<x_{t-1}$.
By Lemma 13, there is a probability of at least $1-\frac{\gamma}{10T}$ that Line 5 well-estimates both intervals $[0,v],[0,w)$ with respect to parameters $\varepsilon/10$ and $\bar{p}=1/100$. Suppose this holds. Then $\hat{\mu}_{\beta_{t}}[0,v]\geq e^{-\varepsilon/10}\mu_{\beta_{t}}[0,v]\geq e^{-1/20}\cdot 1/80>1/90$. So $x_{t}\leq v<x_{t-1}$ and $\mu_{\beta_{t}}[0,x_{t})\leq\mu_{\beta_{t}}[0,v)\leq 1/80$. Likewise, we have $\hat{\mu}_{\beta_{t}}[0,w)\leq e^{1/20}\cdot 1/100<1/90$. Thus $x_{t}\geq w$ and $\mu_{\beta_{t}}[0,x_{t}]\geq\mu_{\beta_{t}}[0,w]\geq 1/100$. ∎
Proposition 24.
If the bounds of Proposition 23 hold, the algorithm does not terminate with ERROR.
Proof.
Suppose the algorithm reaches Line 12, and we have $\beta_{1},\dots,\beta_{T}>\beta_{\min}$ with a decreasing sequence $x_{1}>x_{2}>\dots>x_{T-1}>x_{T}$. For each iteration $t<T$, Proposition 23 gives $\mu_{\beta_{t+1}}[x_{t+1},x_{t})=\mu_{\beta_{t+1}}[0,x_{t})-\mu_{\beta_{t+1}}[0,x_{t+1})\geq 1/4-1/80\geq 1/5$ and $\mu_{\beta_{t}}[x_{t+1},x_{t})\leq\mu_{\beta_{t}}[0,x_{t})\leq 1/80$ and $\mu_{\beta_{t+1}}[0,x_{t})\geq 1/4$. Since the map $\beta\mapsto\mu_{\beta}[0,\chi)$ is non-increasing and $\mu_{\beta_{t+1}}[0,x_{t})\geq 1/4>1/80\geq\mu_{\beta_{t}}[0,x_{t})$, we have $\beta_{t+1}<\beta_{t}$ strictly. We can estimate the partition ratio between $\beta_{t+1}$ and $\beta_{t}$ as:
$\frac{Z(\beta_{t})}{Z(\beta_{t+1})}=\frac{\mu_{\beta_{t+1}}[x_{t+1},x_{t})}{\mu_{\beta_{t}}[x_{t+1},x_{t})}\cdot\frac{\sum_{x\in[x_{t+1},x_{t})}c_{x}e^{\beta_{t}x}}{\sum_{x\in[x_{t+1},x_{t})}c_{x}e^{\beta_{t+1}x}}\geq\frac{1/5}{1/80}e^{(\beta_{t}-\beta_{t+1})x_{t+1}}$ (7)
where the last inequality holds since $x_{t+1}$ is the smallest element in $[x_{t+1},x_{t})$. Alternatively, we can estimate:
$\frac{Z(\beta_{t})}{Z(\beta_{t+1})}=\frac{\mu_{\beta_{t+1}}[0,x_{t-1})}{\mu_{\beta_{t}}[0,x_{t-1})}\cdot\frac{\sum_{x<x_{t-1}}c_{x}e^{\beta_{t}x}}{\sum_{x<x_{t-1}}c_{x}e^{\beta_{t+1}x}}\leq\frac{1}{1/4}e^{(\beta_{t}-\beta_{t+1})x_{t-1}}$ (8)
where again the last inequality holds since every element in $[0,x_{t-1})$ is smaller than $x_{t-1}$.
Inequalities (7) and (8) together show $16e^{(\beta_{t}-\beta_{t+1})x_{t+1}}\leq 4e^{(\beta_{t}-\beta_{t+1})x_{t-1}}$, i.e. $\beta_{t}-\beta_{t+1}\geq\frac{\log 4}{x_{t-1}-x_{t+1}}$. Substituting into (7) and take logarithm gives the bound
$z(\beta_{t+1},\beta_{t})\geq\log 16+\frac{x_{t+1}\log 4}{x_{t-1}-x_{t+1}}\qquad\text{for all $t=1,\dots,T-1$}$ (9)
Now let $g=T/2-2=\lceil\min\{q/2,\sqrt{q\log n}\}\rceil$, and consider the sequence $b_{i}=x_{T-2i}$ for $i=1,\dots g+1$. From Eq. (9), and only including the subset of terms, we have
$\sum_{t=1}^{T-1}z(\beta_{t+1},\beta_{t})\geq\sum_{i=1}^{g}\bigl{(}\log 16+\frac{b_{i}\log 4}{b_{i+1}-b_{i}}\bigr{)}$ (10)
As we show in Lemma 57 in Appendix E, since $b_{1},\dots,b_{g+1}$ is a strictly-increasing positive sequence, we have $\sum_{i=1}^{g}\frac{b_{i}}{b_{i+1}-b_{i}}\geq\frac{g^{2}}{\log(b_{g+1}/b_{1})}-g/2$. Here $b_{g+1}=x_{2}\leq n$, and, since $x_{T-2}<x_{T-1}$ and $x_{T-1}\in\mathcal{F}$, we have $b_{1}\geq 1$. So overall we have shown that
$\sum_{t=1}^{T-1}z(\beta_{t+1},\beta_{t})\geq g\log 16+\frac{g^{2}\log 4}{\log n}-\frac{g\log 4}{2}>\frac{g^{2}}{\log n}+2g.$
Since $\sum_{t=1}^{T-1}z(\beta_{t+1},\beta_{t})\leq z(\beta_{\min},\beta_{\max})\leq q$, this implies that $g<\sqrt{q\log n}$ and $g<q/2$, which is a contradiction to the definition of $g$. ∎
Proposition 25.
Suppose the bounds of Proposition 23 hold. Then with probability at least $41-\gamma/$ 4, the preconditions of Lemma 14 for estimating $\pi$ hold for all $y\in\mathcal{F}$.
Proof.
Our assumption that the call at Line 1 is good immediately gives property (A1) for all $y$.
Consider some iteration $t$ and $y\in(x_{t},x_{t-1}]$, where we denote $x_{t}=-\infty$ for $\beta_{t}=\beta_{\min}$. For brevity write $a_{y}=\mu_{\beta_{t}}(y),\hat{a}_{y}=\hat{\mu}_{\beta_{t}}(y)$. By Proposition 23, if $\beta_{t}>\beta_{\min}$ we have $\mu_{\beta_{t}}[0,y)\geq\mu_{\beta_{t}}[0,x_{t}]\geq 1/100$. Likewise, if $\beta_{t}<\beta_{\max}$, we have $\mu_{\beta_{t}}[y,n]\geq\mu_{\beta_{t}}[x_{t-1},n]\geq 1/4$. So $\beta_{t}\in{\tt BalancingVals}(y,1/100)$. By Proposition 10 this shows $a_{y}\geq\Delta(y)/100$. Thus, condition (A2) will hold as long as
$|\hat{a}_{y}-a_{y}|\leq\varepsilon/3\cdot(a_{y}+\delta/100)$ (11)
For any $y\in\mathcal{F}$, let $\mathcal{E}(y)$ denote the bad event that (11) fails for $y$; note here that we are not necessarily requiring $y\in(x_{t},x_{t-1}]$. Conditioned on $\beta_{t}$, we have
[…]
The map $x\mapsto(1+v/x)^{-Nv/2}$ has second derivative $\frac{Nv^{2}(v(Nv-2)-4x)\left(\frac{v+x}{x}\right)^{-Nv/2}}{4x^{2}(v+x)^{2}}$ which is nonnegative for $x\leq(Nv-2)v/4\leq\delta(\frac{25}{9}\log(\frac{10T}{\delta\gamma})-\frac{\varepsilon}{600})$. In particular, it is concave-up for $x<\delta$. So for $a_{y}\leq\delta$, we can upper-bound the RHS of (12) by its secant line from 00 to $\delta$, i.e.
$F(Na_{y},Nv)\leq\frac{2a_{y}}{\delta}(1+v/\delta)^{-Nv/2}\leq\frac{2a_{y}}{\delta}(1+\varepsilon/300)^{-\frac{5000}{3\varepsilon}\log\frac{10T}{\delta\gamma}}$
Simple calculations show that $(1+\varepsilon/300)^{-\frac{5000}{3\varepsilon}}\leq 0.004$, so this is at most $\frac{2a_{y}}{\delta}\cdot 0.004^{\log\frac{10T}{\delta\gamma}}\leq\frac{a_{y}\gamma}{4T}$.
[…]
The result follows by a union bound over the $T$ iterations. ∎
Overall, the total failure probability is at most $\gamma/4$ (from Line 1) plus $\gamma/2$ (from Proposition 23) plus $\gamma/4$ (from Proposition 25). This concludes the proof of Theorem 22 and the first part of Theorem 1.
5 Solving ${P_{\tt count}}$ for integer-valued Gibbs distributions
The integer-setting algorithms hinge on a data structure called the covering schedule. Formally, we define a covering schedule to be a sequence of the form
$\mathcal{I}=(\beta_{0},w_{0},k_{1},\beta_{1},w_{1},k_{2},\dots,\beta_{t-1},w_{t-1},k_{t},\beta_{t},w_{t})$
which satisfies the following constraints:
(a) $\beta_{\min}=\beta_{0}<\ldots<\beta_{t}=\beta_{\max}$;
(b) $k_{1},\dots,k_{t}\in\mathcal{H}$;
(c) $k_{1}<k_{2}<\dots<k_{t}$
(d) $w_{i}\in[0,1]$ for $i=0,\dots,t$.
Note that $t\leq n+1$. We say that $\mathcal{I}$ is proper if for all $i=1,\dots,t$ it satisfies
$\mu_{\beta_{i-1}}(k_{i})\geq w_{i-1}\text{ and }\mu_{\beta_{i}}(k_{i})\geq w_{i}.$
We define the inverse weight of $\mathcal{I}$, denoted ${\tt InvWt}(\mathcal{I})$ by
${\tt InvWt}(\mathcal{I})=\sum_{i=0}^{t}\frac{1}{w_{i}}.$
Our algorithm to solve ${P^{\delta,\varepsilon}_{\tt count}}$ has three stages. As a high-level summary, it proceeds as follows:
  1. 1.
    Construct a suitable covering schedule $\mathcal{I}=(\beta_{0},w_{0},k_{1},\dots,k_{t},\beta_{t},w_{t})$.
  2. 2.
    Estimate the values $Q(\beta_{i})$ for $i=0,\dots,t$.
  3. 3.
    Use these estimates $\hat{Q}(\beta_{i})$ to estimate the counts $c_{i}$.
Using Theorem 6, we can automatically use this algorithm to solve ${P^{{\varepsilon}}_{\tt ratio-all}}$ as well.
The first stage is quite involved, so we defer it to Section 6 where we show the following result:
Theorem 26.
Procedure ${\tt FindCoveringSchedule}(\gamma)$ produces a covering schedule $\mathcal{I}$, which is proper with probability at least $1-\gamma$.
In the general integer setting, its cost is $O(n(\log^{3}n+\log n\log\tfrac{1}{\gamma}+\log q))$ and ${\tt InvWt}({\cal I})\leq O(n\log n)$.
In the log-concave setting, its cost is $O(n(\log^{2}n+\log\tfrac{1}{\gamma}+\log q))$ and ${\tt InvWt}({\cal I})\leq O(n)$.
For the second stage, we use the following algorithm:
1 if $n\cdot{\tt InvWt(\mathcal{I})}<q\log n$ then
2       for $i=1,\dots t$ form random vars $X_{i}\sim\text{Bernoulli}(\mu_{\beta_{i-1}}(k_{i}))$ and $Y_{i}\sim\text{Bernoulli}(\mu_{\beta_{i}}(k_{i}))$
3       set $\hat{X}^{\tt prod}\leftarrow{\tt EstimateProducts}(X,{\tt InvWt(\mathcal{I})}\cdot 4/\varepsilon^{2},\gamma/2)$
4       set $\hat{Y}^{\tt prod}\leftarrow{\tt EstimateProducts}(Y,{\tt InvWt(\mathcal{I})}\cdot 4/\varepsilon^{2},\gamma/2)$
5       for $i=0,\dots,t$ do set $\hat{Q}(\beta_{i})=\exp\bigl{(}\sum_{j=1}^{i}(\beta_{j}-\beta_{j-1})k_{j}\bigr{)}\cdot\hat{X}^{\tt prod}_{i}/\hat{Y}^{\tt prod}_{i}$
6
7else
8       call $\mathcal{D}\leftarrow{\tt PratioContinuous}(\varepsilon,\gamma)$
9       for $i=0,\dots,t$ do set $\hat{Q}(\beta_{i})=\hat{Q}(\beta_{i}\mid\mathcal{D})$ for all $i=0,\dots,t$
Algorithm 7 Algorithm ${\tt PratioCoveringSchedule}(\mathcal{I},\varepsilon,\gamma)$ for a covering schedule $\mathcal{I}$.
Note that, somewhat counterintuitively, it can be more efficient to use the continuous-setting algorithm to solve ${P^{{\varepsilon}}_{\tt ratio-all}}$, as opposed to directly estimating the values $\hat{Q}(\beta_{i})$ from the covering schedule itself.
Theorem 27.
The algorithm ${\tt PratioCoveringSchedule}(\mathcal{I},\varepsilon,\gamma)$ has cost is $O\bigl{(}\frac{\min\{nW,q\log n,q^{2}\}}{\varepsilon^{2}}\log\tfrac{1}{\gamma}\bigr{)}$ for $W={\tt InvWt}(\mathcal{I})$. If $\mathcal{I}$ is proper, then with probability at least $1-\gamma$ it satisfies
$|\log\hat{Q}(\beta_{i})-\log Q(\beta_{i})|\leq\varepsilon\qquad\qquad\text{forall $i=0,\dots,t$.}$
(When this holds, we say that the call to PratioCoveringSchedule is good).
Proof.
The cost and correctness are clear where $nW\geq q\log n$. So suppose $nW<q\log n$.
Since $t\leq n+1$, Lines $3$ and $4$ have cost $O(nW\varepsilon^{-2}\log\tfrac{1}{\gamma})$. Assuming $\mathcal{I}$ is proper, we have ${{\mathbb{S}}[{X_{i}}]}=\frac{1}{\mu_{\beta_{i-1}}(k_{i})}-1\leq\frac{1}{w_{i-1}}$ for all $i$, so $\sum_{i}{{\mathbb{S}}[{X_{i}}]}\leq W$. Likewise $\sum_{i}{{\mathbb{S}}[{Y_{i}}]}\leq W$. So with probability at least $1-\gamma$, all estimates $\hat{X}^{\tt prod}_{i},\hat{Y}^{\tt prod}_{i}$ are within $e^{\pm\varepsilon/2}$ of $\prod_{j=1}^{i}{\mathbb{E}}[X_{j}],\prod_{j=1}^{i}{\mathbb{E}}[Y_{j}]$ respectively. Observe that
$\frac{{\mathbb{E}}[\prod_{j=1}^{i}X_{j}]}{{\mathbb{E}}[\prod_{j=1}^{i}Y_{j}]}=\prod_{j=1}^{i-1}\frac{\mu_{\beta_{j-1}}(k_{j})}{\mu_{\beta_{j}}(k_{j})}=\prod_{j}e^{(\beta_{j-1}-\beta_{j})k_{j}}\frac{Z(\beta_{j})}{Z(\beta_{j-1})}=Q(\beta_{i})\exp\Bigl{(}\sum_{j=1}^{i}(\beta_{j-1}-\beta_{j})k_{j}\Bigr{)}$
so in that case, the values $\hat{Q}(\beta_{i})$ are also within $e^{\pm\varepsilon}$ of $Q(\beta_{i})$ as required. ∎
5.1 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the general integer setting
We now move on to the third stage, of using the covering schedule to solve ${P_{\tt count}}$. There are two quite distinct algorithms: one for generic integer-valued distributions, and a specialized algorithm for log-concave distributions. We begin with the algorithm for general integer distributions:
1 set $\mathcal{I}=(\beta_{0},w_{0},k_{1},\dots,k_{t},\beta_{t},w_{t})\leftarrow{\tt FindCoveringSchedule}(\gamma/4)$
2 set $(\hat{Q}(\beta_{0}),\dots,\hat{Q}(\beta_{t}))\leftarrow{\tt PratioCoveringSchedule}(\mathcal{I},\varepsilon/9,\gamma/4)$
3 for $i=0,\dots,t$ do let $\hat{\mu}_{\beta_{i}}\leftarrow{\tt Sample}(\beta_{i};\varepsilon/9,w_{i},\frac{\gamma}{10(n+1)^{2}})$
4 for $j=0,\dots,n$ do
5       set $\alpha\leftarrow{\tt Balance}(j,1/4,\frac{\gamma}{4(n+1)^{2}})$
6       find index $i<t$ with $\alpha\in[\beta_{i},\beta_{i+1}]$
7       let $\hat{\mu}_{\alpha}\leftarrow{\tt Sample}(\alpha;\varepsilon/9,\delta/4,\frac{\gamma}{4(n+1)^{2}})$
8       if $\hat{\mu}_{\alpha}(k_{i+1})\geq\delta/4$ then set $\hat{\pi}(j)=\frac{\hat{\mu}_{\beta_{i}}(k_{i+1})}{\hat{\mu}_{\alpha}(k_{i+1})}e^{(\alpha-\beta_{i})k_{i+1}}\hat{Q}(\beta_{i})\cdot e^{-\alpha j}\hat{\mu}_{\alpha}(j)$
9       else if $j<k_{i+1}$    then set $\hat{\pi}(j)=\hat{Q}(\beta_{i})e^{-\beta_{i}j}\hat{\mu}_{\beta_{i}}(j)$
10       else if $j\geq k_{i+1}$   then set $\hat{\pi}(j)=\hat{Q}(\beta_{i+1})e^{-\beta_{i+1}j}\hat{\mu}_{\beta_{i+1}}(j)$
Algorithm 8 Solving problem ${P^{\delta,\varepsilon}_{\tt count}}$.
Theorem 28.
In the integer setting, Algorithm 8 solves ${P^{\delta,\varepsilon}_{\tt count}}$ with cost
$O\Bigl{(}\frac{(n/\delta)\log\tfrac{n}{\gamma}+n^{2}\log n\log\tfrac{1}{\gamma}}{\varepsilon^{2}}+n\log q\Bigr{)}$
Proof.
Lines 1 and 2 have cost $O(\frac{n^{2}\log n\log(1/\gamma)}{\varepsilon^{2}}+n\log q)$. The sampling for $\beta_{i}$ in Line 3 has cost $O(\frac{\log(n/\gamma)}{w_{i}\varepsilon^{2}})$; summing over $i$ gives $O(\frac{{\tt InvWt}(\mathcal{I})}{\varepsilon^{2}}\log\tfrac{n}{\gamma})\leq O(\frac{n\log n}{\varepsilon^{2}}\log\tfrac{n}{\gamma})$. Line 5 has cost $O(n\log\tfrac{nq}{\gamma})$. Line 7 has cost $O(\frac{n}{\delta\varepsilon^{2}}\log\tfrac{n}{\gamma})$. This shows the bound on complexity.
For correctness, assume $\mathcal{I}$ is proper and the calls to Balance and PratioCoveringSchedule are good. Also, assume that Lines 3 and 7 well-estimate every integer in $\mathcal{H}$. By specification of subroutines, these events all hold with probability at least $1-\gamma$. In this case, we claim that the preconditions of Lemma 14 hold for any given $j\in\mathcal{H}$. For this value $j$, let $\alpha,i$ denote the corresponding parameters at Lines 5 and 6 respectively, and define $k=k_{i+1},\beta=\beta_{i}$ for brevity.
By Proposition 10, we have $\mu_{\alpha}(j)\geq\Delta(j)/4$. There are two cases:
  • Suppose $\hat{\mu}_{\alpha}(j)\geq\delta/4$. For (A2), Line 7 well-estimates $j$, so
    $|\hat{\mu}_{\alpha}(j)-\mu_{\alpha}(j)|\leq\varepsilon/9\cdot(\mu_{\alpha}(j)+\delta/4)\leq\varepsilon/9\cdot\mu_{\alpha}(j)(1+\delta/\Delta(j)).$
    For (A1), observe that $Q(\alpha)=\frac{\mu_{\beta}(k)}{\mu_{\alpha}(k)}e^{(\alpha-\beta)k}Q(\beta)$. Line 3 well-estimates $j$ and $\mu_{\beta}(j)\geq w_{i}$, so $|\log\hat{\mu}_{\beta}(j)-\log\mu_{\beta}(j)|\leq\varepsilon/9$. Line 7 well-estimates $k$ and $\hat{\mu}_{\alpha}(k)\geq\delta/4$, so $|\log\hat{\mu}_{\alpha}(k)-\log\mu_{\alpha}(k)|\leq\varepsilon/9$. Also, $|\log\hat{Q}(\beta)-\log Q(\beta)|\leq\varepsilon/9$. Overall $|\log\hat{Q}(\alpha)-\log Q(\alpha)|\leq\varepsilon/3$ where $\hat{Q}(\alpha)=\frac{\hat{\mu}_{\beta}(k_{i+1})}{\hat{\mu}_{\alpha}(k_{i+1})}e^{(\alpha-\beta)k_{i+1}}\hat{Q}(\beta)$.
  • Suppose $\hat{\mu}_{\alpha}(j)<\delta/4$. Let us assume $j<k$; the case with $j\geq k$ is completely analogous. (A1) holds since $|\log\hat{Q}(\beta)-\log Q(\beta)|\leq\varepsilon/9$. Since Line 7 well-estimates $k$ and $\hat{\mu}_{\alpha}(k)<\delta/4$ we have $\mu_{\alpha}(k)<e^{\varepsilon/9}\cdot\delta/4\leq\delta/2$. We have already seen that $\mu_{\alpha}(j)\geq\Delta(j)/4$ and since $\mathcal{I}$ is proper, we have $\mu_{\beta}(k)\geq w_{i}$. So from the Uncrossing Inequality we get $\mu_{\beta}(j)\geq\frac{\mu_{\beta}(k)\mu_{\alpha}(j)}{\mu_{\alpha}(k)}\geq\frac{w_{i}\Delta(j)/4}{\delta/2}=\frac{w_{i}\Delta(j)}{2\delta}.$ Now Line 7 well-estimates $j$ so
    $|\hat{\mu}_{\beta}(j)-\mu_{\beta}(j)|\leq\varepsilon/9\cdot(\mu_{\beta}(j)+w_{i})\leq\varepsilon/9\cdot(\mu_{\beta}(j)+\frac{2\delta\mu_{\beta}(j)}{\Delta(j)})\leq 2\varepsilon/9\cdot\mu_{\beta}(j)(1+\frac{\delta}{\Delta(j)}).\qed$
With some simplification of parameters, this gives the second part of Theorem 1; via Theorem 6, it also gives the second part of Theorem 3.
5.2 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the log-concave setting
As we have mentioned, there is an alternative algorithm to estimate counts in the log-concave setting:
1 set $\mathcal{I}=(\beta_{0},w_{0},k_{1},\dots,k_{t},\beta_{t},w_{t})\leftarrow{\tt FindCoveringSchedule}(\gamma/3)$
2 set $(\hat{Q}(\beta_{0}),\dots,\hat{Q}(\beta_{t}))\leftarrow{\tt PratioCoveringSchedule}(\mathcal{I},\varepsilon/3,\gamma/3)$
3 set $w_{0}\leftarrow\min\{w_{0},\delta,1/n\}$ and set $w_{t}\leftarrow\min\{w_{t},\delta,1/n\}$
4 for $i=0,\dots,t$ do let $\hat{\mu}_{\beta_{i}}\leftarrow{\tt Sample}(\beta_{i};\varepsilon/9,w_{i},\frac{\gamma}{3(n+1)^{2}})$
5 for $i=1,\dots t-1$ do
6       for $j=k_{i}+1,\dots,k_{i+1}$ do set $\hat{\pi}(j)=\hat{Q}(\beta_{i})e^{-\beta_{i}j}\hat{\mu}_{\beta_{i}}(j)$
7for $j=0,\dots,k_{1}$ do                 set $\hat{\pi}(j)=\hat{Q}(\beta_{\min})e^{-\beta_{\min}j}\hat{\mu}_{\beta_{\min}}(j)$
for $j=k_{t}+1,\dots,n$ do           set $\hat{\pi}(j)=\hat{Q}(\beta_{\max})e^{-\beta_{\max}j}\hat{\mu}_{\beta_{\max}}(j)$
Algorithm 9 Solving ${P^{\delta,\varepsilon}_{\tt count}}$ in the log-concave setting.
Theorem 29.
In the log-concave setting, Algorithm 9 solves ${P^{\delta,\varepsilon}_{\tt count}}$ with cost
$O\Bigl{(}n\log^{2}n+n\log q+\frac{\min\{n^{2},q\log n\}\log\frac{1}{\gamma}+(n+1/\delta)\log\frac{n}{\gamma}}{\varepsilon^{2}}\Bigr{)}$
Proof.
Note that ${\tt InvWt}(\mathcal{I})\leq O(n)$. So Line 1 has cost $O(n(\log^{2}n+\log q+\log\frac{1}{\gamma}))$, and Line 2 has cost $O(\frac{\min\{n^{2},q\log n\}}{\varepsilon^{2}}\log\frac{1}{\gamma})$. Because of the modification step at Line 3, the total cost of Line 4 is $O\bigl{(}(\frac{1}{\delta}+n+\sum_{i=0}^{t}\frac{1}{w_{i}})\varepsilon^{-2}\log\tfrac{n}{\gamma}\bigr{)}=O(\frac{(n+1/\delta)}{\varepsilon^{2}}\log\frac{n}{\gamma})$. These add to the stated complexity.
For correctness, suppose that $\mathcal{I}$ is proper and the call to PratioCoveringSchedule is good and every value $j\in\mathcal{H}$ is well-estimated by every iteration of Line 4. By specification of the subroutines, these events hold with probability at least $1-\gamma$. We claim that the preconditions of Lemma 14 then hold for all $j\in\mathcal{H}$. There are three cases to consider.
  • Suppose $j\in(k_{i},k_{i+1}]$ is estimated at Line 6. Let $a=\mu_{\beta_{i}}(j),\hat{a}=\hat{\mu}_{\beta_{i}}(j)$. Condition (A1) holds since $|\log\hat{Q}(\beta_{i})-\log Q(\beta_{i})|\leq\varepsilon/3$. Note that $a\geq\min\{\mu_{\beta_{i}}(k_{i}),\mu_{\beta_{i}}(k_{i+1})\}\geq w_{i}$, where the first inequality follows from log-concavity of counts and the second inequality holds from properness of $\mathcal{I}$. Since Line 4 well-estimates $j$, this implies $|a-\hat{a}|\leq\varepsilon/6\cdot(a+w_{i})\leq\varepsilon/3\cdot a$, giving (A2).
  • Suppose $j\leq k_{1}$ is estimated at Line 7, and let $a=\mu_{\beta_{\min}}(j),\hat{a}=\hat{\mu}_{\beta_{\min}}(j)$. Again, condition (A1) is immediate. Since Line 4 well-estimates $j$, we have $|\hat{a}-a|\leq\varepsilon/6\cdot(a+w_{0}^{\prime})$ where $w_{0}^{\prime}=\min\{w_{0},\delta,1/n\}$ is the modified value and $w_{0}$ is the original value from $\mathcal{I}$. To show (A2), it thus suffices to show that $a\geq w_{0}^{\prime}$.
    If $j$ is on the “decreasing” slope of the log-concave distribution $\mu_{\beta_{\min}}$, then properness of $\mathcal{I}$ implies that $a\geq\mu_{\beta_{\min}}(k_{1})\geq w_{0}$ which immediately gives (A2). So suppose $j$ is on the increasing slope of $\mu_{\beta_{\min}}$, and hence $\mu_{\beta_{\min}}(\ell)\leq\mu_{\beta_{\min}}(j)=a$ for all $\ell\leq j$. So $\mu_{\beta_{\min}}[0,j-1]\leq ja$.
    Let $\Delta(j)=\mu_{\alpha}(j)$ for $\alpha\in[\beta_{\min},\beta_{\max}]$. For $\ell\geq j$, the Uncrossing Inequality gives $\mu_{\beta_{\min}}(\ell)\leq\frac{\mu_{\alpha}(\ell)\mu_{\beta_{\min}}(j)}{\mu_{\alpha}(j)}=\frac{\mu_{\alpha}(\ell)a}{\Delta(j)}$, so $\mu_{\beta_{\min}}[j,n]\leq\frac{a}{\Delta(j)}\cdot\mu_{\alpha}[j,n]\leq\frac{a}{\Delta(j)}$. So $1=\mu_{\beta_{\min}}[0,j-1]+\mu_{\beta_{\min}}[j,n]\leq ja+\frac{a}{\Delta(j)}$, implying $a\geq\frac{\Delta(j)}{1+j\Delta(j)}$.
    So $a(1+\delta/\Delta(j))\geq\frac{\delta+\Delta(j)}{1+j\Delta(j)}\geq\min\{\delta,1/j\}\geq w_{0}^{\prime}$. This establishes (A2).
  • Suppose $j>k_{t+1}$ is estimated at Line 8. This is completely symmetric to the previous case. ∎
Again, with some simplification of parameters, this gives the third part of Theorem 1; via Theorem 6, it also gives the third part of Theorem 3.
6 Constructing a covering schedule
We will show the following more precise bound on the weight of the schedule.
Theorem 30.
Let $a>4$ be an arbitrary constant and define the following parameter [^3] As motivation for the definition of $\rho$, see Lemma 58.
$\rho{\stackrel{{\scriptstyle\mbox{\tiny\tt~{}def~{}}}}{{=}}}\begin{cases}1+\log(n+1)&\mbox{in the general integer setting}\\ e&\mbox{in the log-concave setting}\end{cases}$
In either the general integer or log-concave setting, the procedure ${\tt FindCoveringSchedule}(\gamma)$ produces a covering schedule $\mathcal{I}$ with ${\tt InvWt}({\cal I})\leq a(n+1)\rho$, which is proper with probability at least $1-\gamma$. It has cost $O(n\rho(\log^{2}n+\log\frac{1}{\gamma})+n\log q)$.
This will immediately imply Theorem 26. The construction has two parts. First, in Section 6.1, we build an object with relaxed constraints called a preschedule. Then in Section 6.2, we convert this into a schedule.
6.1 Constructing a preschedule
For Section 6, let us fix constants $\tau\in(0,\frac{1}{2})$, $\lambda>1$, and define parameter $\phi=\frac{\tau}{\lambda^{3}\rho}$. Thus, $\phi=\Theta(\frac{1}{\log n})$ in the general setting and $\phi=\Theta(1)$ in the log-concave setting. We introduce basic terminology and definitions.
An $\mathcal{H}$-interval is a discrete set of points $\{\sigma^{-},\sigma^{-}+1,\dots,\sigma^{+}-1,\sigma^{+}\}$, for integers $0\leq\sigma^{-}\leq\sigma^{+}\leq n$. We also write this more compactly as $\sigma=[\sigma^{-},\sigma^{+}]$; note that the set $\sigma$ has cardinality $|\sigma|=\sigma^{+}-\sigma^{-}+1$.
A segment is a tuple $\theta=(\beta,\sigma)$ where $\beta\in[\beta_{\min},\beta_{\max}]$, and $\sigma$ is an $\mathcal{H}$-interval. We say $\theta$ is $\phi$-proper (or just proper if $\phi$ is understood) if it satisfies the following two properties:
  • Either $\beta=\beta_{\min}$ or $\mu_{\beta}(\sigma^{-})\geq\phi/|\sigma|$
  • Either $\beta=\beta_{\max}$ or $\mu_{\beta}(\sigma^{+})\geq\phi/|\sigma|$
A preschedule is a sequence of segments ${\cal J}=((\beta_{0},\sigma_{0}),\ldots,(\beta_{t},\sigma_{t}))$ with the following properties:
  • (I0)
    $\sigma_{i+1}^{-}\leq\sigma_{i}^{+}$ for $i=0,\dots,t-1$.
  • (I1)
    $\beta_{\min}=\beta_{0}\leq\ldots\leq\beta_{t}=\beta_{\max}$.
  • (I2)
    $0=\sigma^{-}_{0}\leq\ldots\leq\sigma^{-}_{t}\leq n$ and $0\leq\sigma^{+}_{0}\leq\ldots\leq\sigma^{+}_{t}=n$
We say that $\cal J$ is $\phi$-proper if all segments $\theta_{i}$ are $\phi$-proper. The main idea of the algorithm is to maintain a sequence of proper segments satisfying properties (I1) and (I2), and grow it until it satisfies (I0). Details are provided below:
1 call $\theta_{\min}\,\leftarrow{\tt FindSegment}(\beta_{\min},\,[0,0],[0,n])$ and $\theta_{\max}\leftarrow{\tt FindSegment}(\beta_{\max},[0,n],[n,n])$
2 initialize $\mathcal{J}$ to contain the two segments $\theta_{\min},\theta_{\max}$
3 while $\mathcal{J}$ does not satisfy (I0) do
4       pick arbitrary consecutive segments $\theta_{{\tt left}}=(\beta_{{\tt left}},\sigma_{{\tt left}})$ and $\theta_{{\tt right}}=(\beta_{{\tt right}},\sigma_{{\tt right}})$ in $\mathcal{J}$ with $\sigma_{{\tt left}}^{+}<\sigma_{{\tt right}}^{-}$.
5       let $M=\big{\lfloor}\frac{\sigma_{{\tt left}}^{+}+\sigma_{{\tt right}}^{-}}{2}\big{\rfloor}+\frac{1}{2}$.
6       call $\beta\leftarrow{\tt Balance}(M,\tau,\frac{1}{4n})$
7       call $\theta\leftarrow\begin{cases}{\tt FindSegment}(\beta,\ \ \ \ \ \thinspace[\sigma^{-}_{{\tt left}},M-\tfrac{1}{2}],\thinspace[M+\tfrac{1}{2},\sigma^{+}_{{\tt right}}])&\mbox{if }\beta_{{\tt left}}<\beta<\beta_{{\tt right}}\\ {\tt FindSegment}(\beta_{{\tt left}},\ [\sigma^{-}_{{\tt left}},\sigma^{-}_{{\tt left}}],\ \ \thinspace[M+\tfrac{1}{2},\sigma^{+}_{{\tt right}}])&\mbox{if }\beta\leq\beta_{{\tt left}}\\ {\tt FindSegment}(\beta_{{\tt right}},[\sigma^{-}_{{\tt left}},M-\tfrac{1}{2}],[\sigma^{+}_{{\tt right}},\sigma^{+}_{{\tt right}}])&\mbox{if }\beta\geq\beta_{{\tt right}}\end{cases}$
8       insert $\theta$ into $\mathcal{J}$ between $\theta_{{\tt left}}$ and $\theta_{{\tt right}}$
return $\mathcal{J}$     Subroutine ${\tt FindSegment}(\beta,[h_{{\tt left}},a_{{\tt left}}],[a_{{\tt right}},h_{{\tt right}}])$:
9 let $\hat{\mu}_{\beta}\leftarrow{\tt Sample}(\beta;\tfrac{1}{2}\log\lambda,\bar{p},\frac{1}{4(n+2)^{2}})$ where $\bar{p}=\frac{\phi}{h_{{\tt right}}-h_{{\tt left}}+1}$.
10 for $i=h_{{\tt left}}$ to $h_{{\tt right}}$ set $\Phi(i)=\begin{cases}\hat{\mu}_{\beta}(i)&\mbox{if }i\notin\{h_{{\tt left}},h_{{\tt right}}\}\\ \lambda\cdot\hat{\mu}_{\beta}(i)&\mbox{if }i\in\{h_{{\tt left}},h_{{\tt right}}\}\\ \end{cases}\vspace{0.05in}$
11 set $k^{-}=\negthickspace\displaystyle\operatorname*{arg\,max}_{i\in[h_{{\tt left}},a_{{\tt left}}]}(a_{{\tt left}}-i+1)\Phi(i)$ and $k^{+}=\negthickspace\displaystyle\operatorname*{arg\,max}_{i\in[a_{{\tt right}},h_{{\tt right}}]}(i-a_{{\tt right}}+1)\Phi(i)$
return segment $\theta=(\beta,\sigma)$ where $\sigma=[k^{-},k^{+}]$
Algorithm 10 ${\tt FindSegment}(\beta,[h_{{\tt left}},a_{{\tt left}}],[a_{{\tt right}},h_{{\tt right}}])$.
Algorithm 11 Computing an initial preschedule.
At each iteration, the uncovered set $[0,n]-\bigcup_{(\beta,\sigma)\in\mathcal{J}}\sigma$ is a union of disjoint open intervals (“gaps”). The interval $L=(\sigma^{+}_{{{\tt left}}},\sigma^{-}_{{{\tt right}}})$ at an iteration is one of these gaps. From the new segment produced by FindSegment, this gap $L$ gets replaced at the next iteration by two new gaps $L^{\prime},L^{\prime\prime}$. Overall, the set of gaps over all the iterations forms a laminar family of intervals. Between gaps, the “filled” regions have strictly increasing values of $\beta$. Thus at Line 4 we have $\beta_{\tt left}<\beta_{\tt right}$ strictly.
The loop in Lines 3 – 8 is executed at most $n$ times, since each time it covers a new half-integer value $M$. So there are at most $n+2$ calls to ${\tt FindSegment}$ and at most $n$ calls to ${\tt Balance}$.
Proposition 31.
Algorithm 11 has cost $O(n\log q+n\rho\log^{2}n)$.
Proof.
By Theorem 11, the ${\tt Balance}$ subroutines have cost $O(n\log(nq))$. The two calls to FindSegment at Line 1 have cost $O(\rho n\log n)$. Let ${\cal J}_{i}$ be the sequence and $M_{i},\sigma_{{{\tt left}},i},\sigma_{{{\tt right}},i}$ be the variables at the $i^{\text{th}}$ iteration and let $A_{i}=[\sigma_{{{\tt left}},i}^{-},\sigma_{{{\tt right}},i}^{+}]$ and $L_{i}=(\sigma_{{\tt left},i}^{+},\sigma_{{\tt right},i}^{-})$. We can observe that the call to ${\tt FindSegment}$ at the $i^{\text{th}}$ iteration has cost $O(|A_{i}|\rho\log n)$, noting that $\sigma_{{\tt right},i}^{+}>\sigma_{{\tt left},i}^{-}$ strictly. We now show that $\sum_{i}|A_{i}|=O(n\log n)$, which will yield the claim about the complexity.
For each $k\in{\cal H}$ define $I^{-}_{k}=\{i\>:\>k\in A_{i}\;\wedge\;M_{i}<k\}$, and consider $i,j\in I^{-}_{k}$ with $i<j$. We claim that intervals $L_{i},L_{j}$ overlap. For, suppose $L_{j}$ comes strictly after $L_{i}$. Since $L_{j}$ is always selected from a gap at iteration $j$, it cannot intersect with $\sigma_{{\tt right},i}$, and in particular $\sigma_{{\tt left},j}^{+}\geq\sigma_{{\tt right},i}^{+}$. In this case, by definition of $M_{j}$ and $I_{k}$, we would have $k>M_{j}>\sigma_{{\tt left},j}^{+}$ and $k\leq\sigma_{{\tt right},i}^{+}$, a contradiction.
On the other hand, suppose $L_{j}$ comes strictly before $L_{i}$. Again, since $L_{j}$ is selected from a gap, it comes strictly before $\sigma_{{\tt left},i}$ and so $\sigma_{{\tt right},j}^{-}\leq\sigma_{{\tt left},i}^{-}$. By Property (I2) this implies $\sigma_{{\tt right},j}$ comes before $\sigma_{{\tt left},i}$ and so also $\sigma_{{\tt right},j}^{+}\leq\sigma_{{\tt left},i}^{+}$. Then by definition of $M_{i}$ and $I_{k}^{-}$ we would have $\sigma_{{\tt left},i}^{+}<M_{i}<k$ and $k\leq\sigma_{{\tt right},j}^{+}$, again a contradiction.
So $L_{i}$ and $L_{j}$ overlap. Since $M_{i}$ is chosen near the median of $L_{i}$, we must have $|L_{j}|\leq\tfrac{1}{2}|L_{i}|$. As this holds for all pairs $i,j\in I^{-}_{k}$, we conclude that $|I^{-}_{k}|\leq 1+\log_{2}n$. Similarly, $|I^{+}_{k}|\leq 1+\log_{2}n$ where $I^{+}_{k}=\{i\>:\>k\in A_{i}\;\wedge\;M_{i}>k\}$. It remains to observe that $\sum_{i}|A_{i}|=\sum_{k\in{\cal H}}|I^{-}_{k}\cup I^{+}_{k}|$. ∎
The main result for this theorem, which is technically involved, is the following:
Lemma 32.
With probability at least $1/2$, the final preschedule $\mathcal{J}$ is proper.
In order to show Lemma 32, we will suppose for the remainder of the section that all calls to Balance are good, and that Line 10 in every call to FindSegment well-estimates every value $j\in\mathcal{H}$. By specification of subroutines, these properties hold with property at least $1/2$.
Indeed, we will show by induction on iteration count that the following two additional invariants are preserved in this case:
  • (I3)
    Each segment $\theta$ of $\mathcal{J}$ is $\phi$-proper.
  • (I4)
    Each segment $\theta$ of $\mathcal{J}$ satisfies two “extremality” conditions:
    $\mu_{\beta}(k)$$\leq$ $\lambda\cdot\frac{|\sigma|}{|\sigma|+(\sigma^{-}-k)}\cdot\mu_{\beta}(\sigma^{-})\qquad\quad\forall k\in\{0,\ldots,\sigma^{-}-1\}$ (13a)
    $\mu_{\beta}(k)$$\leq$ $\lambda\cdot\frac{|\sigma|}{|\sigma|+(k-\sigma^{+})}\cdot\mu_{\beta}(\sigma^{+})\qquad\quad\forall k\in\{\sigma^{+}\!+\!1,\ldots,n\}$ (13b)
Note that FindSegment gives a slight bias to the endpoints $h_{{\tt left}}$ or $h_{{\tt right}}$, in the definition of $\Phi$ at Line 11. This is needed to preserve the slack factor $\lambda>1$ in the definition of extremality (13a),(13b). Without this bias, the factor would grow uncontrollably as the algorithm progresses.
Now we consider some call to FindSegment, made at Line 1 or some iteration of Line 7. By induction, properties (I3) and (I4) hold for all segments in $\mathcal{J}$ up to this point. We need to show that the segment $\theta$ returned by FindSegment also satisfies these properties. The cases when FindSegment is called in Line 1, or in Line 7 when $\beta\in\{\beta_{{\tt left}},\beta_{{\tt right}}\}$, are handled very differently from the main case, which is Line 7 with $\beta\in(\beta_{{\tt left}},\beta_{{\tt right}})$ strictly. In the former cases, there is no “free choice” for the left-margin $k^{-}$ or right-margin $k^{+}$ respectively; for instance, when $\beta=\beta_{{\tt left}}$, then $h_{{\tt left}}=a_{{\tt left}}=\sigma_{{\tt left}}^{-}$ and our only choice is to set $k^{-}=\sigma_{{\tt left}}^{-}$.
We say the call to FindSegment at Line 1 with $\beta=\beta_{\min}$, or the call at Line 7 with $\beta=\beta_{{\tt left}}$, is left-forced; the call at Line 1 with $\beta=\beta_{\max}$, or at Line 7 with $\beta=\beta_{{\tt right}}$ is right-forced. Otherwise it is left-free and right-free respectively.
Lemma 33.
(a) If the call is left-free, then $(a_{{\tt left}}-i+1)\mu_{\beta}(i)\leq\lambda(a_{{\tt left}}-h_{{\tt left}}+1)\mu_{\beta}(h_{{\tt left}})$ for all $i<h_{{\tt left}}$.
(b) If the call is right-free, then $(i-a_{{\tt right}}+1)\mu_{\beta}(i)\leq\lambda(h_{{\tt right}}-a_{{\tt right}}+1)\mu_{\beta}(h_{{\tt right}})$ for all $i>h_{{\tt right}}$.
Proof.
We only show (a); the proof of (b) is analogous. If FindSegment is called at Line 1 then $h_{{\tt left}}=0$ and the claim is vacuous. So assume FindSegment is called at Line 7, and consider $i<h_{{\tt left}}$. Since segment $\theta_{{\tt left}}=(\beta_{{\tt left}},\sigma_{{\tt left}})$ satisfies Eq. (13a) and $h_{{\tt left}}=\sigma_{{\tt left}}^{-}$, we have
$\mu_{\beta_{{\tt left}}}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\cdot\mu_{\beta_{{\tt left}}}(h_{{\tt left}})$ (14)
Since $\beta>\beta_{{\tt left}}$, the Uncrossing Inequality gives $\mu_{\beta_{{\tt left}}}(i)\mu_{\beta}(h_{{\tt left}})\geq\mu_{\beta_{{\tt left}}}(h_{{\tt left}})\mu_{\beta}(i)$. Combined with Eq. (14), this yields
$\mu_{\beta}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\cdot\mu_{\beta}(h_{{\tt left}})$
Finally, since the call is left-free at Line 7, we have $a_{{\tt left}}\geq\sigma_{{\tt left}}^{+}$, so $|\sigma_{{\tt left}}|\leq a_{{\tt left}}+1-h_{{\tt left}}$ and
$\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(h_{{\tt left}}-i)}\leq\frac{a_{{\tt left}}+1-h_{{\tt left}}}{(a_{{\tt left}}+1-h_{{\tt left}})+(h_{{\tt left}}-i)}=\frac{a_{{\tt left}}-h_{{\tt left}}+1}{a_{{\tt left}}-i+1}\qed$
Lemma 34.
(a) If the call is left-free, there is $k\in[h_{{\tt left}},a_{{\tt left}}]$ with $(a_{{\tt left}}-k+1)\cdot\mu_{\beta}(k)\geq\phi\lambda^{2}$.
(b) If the call is right-free, there is $k\in[a_{{\tt right}},h_{{\tt right}}]$ with $(k-a_{{\tt right}}+1)\cdot\mu_{\beta}(k)\geq\phi\lambda^{2}$.
Proof.
The two parts are completely analogous, so we only prove (a). We first observe that
$\mu_{\beta}[0,a_{{\tt left}}]\geq\tau.$ (15)
This is trivial if FindSegment is called at Line 1 with $a_{{\tt left}}=n$ in which case $\mu_{\beta}[0,a_{{\tt left}}]=1$. Otherwise, if FindSegment is called at Line 7 and the call is left-free, we have $\beta\in{\tt BalancingVals}(M,\tau)$ where $M=a_{{\tt left}}+1/2$.
Now assume (a) is false i.e. $(a_{{\tt left}}+1-k)\cdot\mu_{\beta}(k)<\phi\lambda^{2}=\frac{\tau}{\lambda\rho}$ for all $k\in[h_{{\tt left}},a_{{\tt left}}]$. We will use this to derive a contradiction to Eq. (15). Under the assumption that (a) is false, we claim the following:
$(a_{{\tt left}}+1-k)\mu_{\beta}(k)<\tau/\rho\qquad\qquad\text{for all }k\leq a_{{\tt left}}$ (16)
Indeed, we have already assumed the stronger inequality $(a_{{\tt left}}+1-k)\mu_{\beta}(k)<\frac{\tau}{\lambda\rho}$ for $k\in[h_{{\tt left}},a_{{\tt left}}]$. In particular, we know
$(a_{{\tt left}}+1-h_{{\tt left}})\mu_{\beta}(h_{{\tt left}})<\frac{\tau}{\lambda\rho}$ (17)
For $k<h_{{\tt left}}$, Lemma 33(a) with $i=k$ gives $(a_{{\tt left}}+1-k)\mu_{\beta}(k)\leq\lambda(a_{{\tt left}}+1-h_{{\tt left}})\mu_{\beta}(h_{{\tt left}})$. Combined with Eq. (17), this gives the bound of Eq. (16): $(a_{{\tt left}}+1-k)\mu_{\beta}(k)<\lambda\cdot\frac{\tau}{\lambda\rho}=\tau/\rho$.
Now let $\ell=a_{{\tt left}}+1$, and consider the sequence
$b_{i}=\mu_{\beta}(a_{{\tt left}}+1-i)\cdot\rho/\tau\qquad\qquad\text{for $i=1,\dots,\ell$}$
By Eq. (16), we have $b_{i}\leq 1/i$ for all $i$. We claim that $\sum_{i=1}^{\ell}b_{i}<\rho$. There are two cases.
  • Log-concave setting with $\boldsymbol{\rho=e}$. Since the counts are log-concave, so is the sequence $b_{i}$ (since $\mu_{\beta}(k)\propto c_{k}e^{\beta k}$). As we show in Lemma 58 in Appendix E, this implies that $\sum_{i}b_{i}<e$.
  • General setting with $\boldsymbol{\rho=1+\log(n+1)}$. We have $\sum_{i=1}^{\ell}b_{i}<1+\log\ell\leq 1+\log(n+1)$ by the well-known inequality for the harmonic series.
Now observe that $\mu_{\beta}[0,a_{{\tt left}}]=\sum_{i}\frac{\tau}{\rho}\cdot b_{i}<\tau$. This indeed contradicts Eq. (15). ∎
Proposition 35.
(a) If the call is left-free, then $(a_{{\tt left}}-k^{-}+1)\cdot\hat{\mu}_{\beta}(k^{-})\geq\phi\sqrt{\lambda}$ and $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}$.
(b) If the call is right-free, then $(k^{+}-a_{{\tt right}}+1)\cdot\hat{\mu}_{\beta}(k^{+})\geq\phi\sqrt{\lambda}$ and $\mu_{\beta}(k^{+})\geq\hat{\mu}_{\beta}(k^{+})/\sqrt{\lambda}$.
Proof.
We only prove (a); the case (b) is completely analogous.
By Lemma 34, there is $k\in[h_{{\tt left}},a_{{\tt left}}]$ with $\mu_{\beta}(k)\geq\frac{\phi\lambda^{2}}{a_{{\tt left}}-k+1}\geq\bar{p}$. Since Line 10 well-estimates $k$, we have $\Phi(k)\geq\hat{\mu}_{\beta}(k)\geq\mu_{\beta}(k)/\sqrt{\lambda}\geq\frac{\phi\lambda^{3/2}}{a_{{\tt left}}+1-k}$. Since $k^{-}$ is chosen as the argmax, $(a_{{\tt left}}+1-k^{-})\Phi(k^{-})\geq(a_{{\tt left}}+1-k)\Phi(k)\geq\phi\lambda^{3/2}$. So $\hat{\mu}_{\beta}(k^{-})\geq\Phi(k^{-})/\lambda\geq\frac{\phi\sqrt{\lambda}}{a_{{\tt left}}+1-k^{-}}\geq\bar{p}$. Since Line 10 well-estimates $k^{-}$, this implies $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}$. ∎
Proposition 36.
The segment $\theta=(\beta,\sigma)$ produced at the given iteration is $\phi$-proper.
Proof.
We need to show that if $\beta>\beta_{\min}$ then $\mu_{\beta}(k^{-})\geq\frac{\phi}{|\sigma|}$ and likewise if $\beta<\beta_{\max}$ then $\mu_{\beta}(k^{+})\geq\frac{\phi}{|\sigma|}$. We show the former; the latter is completely analogous.
If the call is left-forced, and $\beta\neq\beta_{\min}$, then necessarily $\beta=\beta_{{\tt left}}$ and $k^{-}=\sigma_{{\tt left}}^{-}$. Since $\theta_{{\tt left}}$ satisfies (I3), we have $\mu_{\beta}(k^{-})\geq\frac{\phi}{|\sigma_{{\tt left}}|}\geq\frac{\phi}{|\sigma|}$.
If the call is left-free, then Proposition 35 gives $\mu_{\beta}(k^{-})\geq\hat{\mu}_{\beta}(k^{-})/\sqrt{\lambda}\geq\frac{\phi}{a_{{\tt left}}+1-k^{-}}\geq\frac{\phi}{|\sigma|}$. ∎
Proposition 37.
The segment $\theta=(\beta,\sigma)$ produced at the given iteration satisfies properties (I4).
Proof.
We only verify that $\theta$ satisfies Eq. (13a); the case of Eq. (13b) is completely analogous.
First, suppose the call is left-forced. If $h_{{\tt left}}=0$ there is nothing to show. If FindSegment is called at Line 7 with $\beta=\beta_{{\tt left}}$, then $k^{-}=\sigma_{{\tt left}}^{-},k^{+}\geq a_{{\tt right}}\geq\sigma_{{\tt left}}^{+}$; so $|\sigma|\geq|\sigma_{{\tt left}}|$. For $i<k^{-}$, we use Eq. (13a) for segment $\theta_{{\tt left}}$ to calculate:
$\mu_{\beta}(i)\leq\lambda\cdot\frac{|\sigma_{{\tt left}}|}{|\sigma_{{\tt left}}|+(\sigma_{{\tt left}}^{-}-i)}\cdot\mu_{\beta}(\sigma_{{\tt left}}^{-})\leq\lambda\cdot\frac{|\sigma|}{|\sigma|+(\sigma^{-}-i)}\cdot\mu_{\beta}(k^{-})$
Otherwise, suppose the call is left-free. Since $|\sigma|\geq a_{{\tt left}}-k^{-}+1$, it suffices to show that
$(a_{{\tt left}}-i+1)\mu_{\beta}(i)\leq\lambda(a_{{\tt left}}-k^{-}+1)\mu_{\beta}(k^{-})\qquad\quad\text{for }i<k^{-}$ (18)
If $k^{-}=h_{{\tt left}}$, this is precisely Lemma 33(a). So suppose that $k^{-}>h_{{\tt left}}$ and so $\Phi(k^{-})=\hat{\mu}_{\beta}(k^{-})$. Define $\kappa_{j}=1$ for $j\notin\{h_{{\tt left}},h_{{\tt right}}\}$ and $\kappa_{j}=\lambda$ for $j\in\{h_{{\tt left}},h_{{\tt right}}\}$, i.e. $\Phi(j)=\kappa_{j}\cdot\hat{\mu}_{\beta}(j)$. Now let $i\in\{h_{{\tt left}},\dots,k^{-}\}$. By definition of $k^{-}$, we have $(a_{{\tt left}}-i+1)\Phi(i)\leq(a_{{\tt left}}-k^{-}+1)\Phi(k^{-})$, i.e.
$\hat{\mu}_{\beta}(i)\leq\frac{(a_{{\tt left}}-k^{-}+1)\Phi(k^{-})}{\kappa_{i}(a_{{\tt left}}-i+1)}=\frac{(a_{{\tt left}}-k^{-}+1)\hat{\mu}_{\beta}(k^{-})}{\kappa_{i}(a_{{\tt left}}-i+1)}$ (19)
By Proposition 35, $(a_{{\tt left}}-k^{-}+1)\hat{\mu}_{\beta}(k^{-})\geq\phi\sqrt{\lambda}$. In particular, the RHS of Eq. (19) is at least $\bar{p}$. Since Line 1 well-estimates $i$ and since $\hat{\mu}_{\beta}(k^{-})\leq\sqrt{\lambda}\mu_{\beta}(k^{-})$, this implies
$\mu_{\beta}(i)\leq\frac{\lambda(a_{{\tt left}}-k^{-}+1)\mu_{\beta}(k^{-})}{\kappa_{i}(a_{{\tt left}}-i+1)}.$ (20)
For $i\in\{h_{{\tt left}}+1,\dots,k^{-}\}$, we have $\kappa_{i}=1$, and so Eq. (20) is exactly Eq. (18). For $i=h_{{\tt left}}$, we have $\kappa_{i}=\lambda$ and so Eq. (20) shows
$(a_{{\tt left}}-h_{{\tt left}}+1)\mu_{\beta}(h_{{\tt left}})\leq(a_{{\tt left}}-k^{-}+1)\mu_{\beta}(k^{-})$ (21)
which again establishes Eq. (18) (with additional slack). Finally, when $i<h_{{\tt left}}$, we show Eq. (18) by combining Lemma 33(a) with Eq. (21):
$(a_{{\tt left}}-i+1)\mu_{\beta}(i)\leq\lambda(a_{{\tt left}}-h_{{\tt left}}+1)\mu_{\beta}(h_{{\tt left}})\leq\lambda(a_{{\tt left}}-k^{-}+1)\mu_{\beta}(k^{-})\qed$
Thus, the preschedule $\mathcal{J}$ at the end maintains property (I3), and hence it is $\phi$-proper. This concludes the proof of Lemma 32.
6.2 Converting the preschedule into a covering schedule
There are two steps to convert the preschedule into a covering schedule. First, we throw away redundant intervals; second, we “uncross” the adjacent intervals. While we are doing this, we also check if the resulting schedule is proper; if not, we will discard it and generate a new preschedule from scratch. Here $\nu>0$ is some arbitrary constant.
1 while $\mathcal{J}$ contains any adjacent segments $\theta_{i}=(\beta_{i},\sigma_{i}),\theta_{i+1}=(\beta_{i+1},\sigma_{i+1})$ with $\beta_{i}=\beta_{i+1}$ do
2       merge the segments, i.e. replace $\theta_{i},\theta_{i+1}$ with a single segment $\theta^{\prime}=(\beta_{i},[\sigma_{i}^{-},\sigma_{i+1}^{+}])$.
3while $\mathcal{J}$ contains any index $i\in\{1,\dots,t-1\}$ with $\sigma_{i+1}^{-}\leq\sigma_{i-1}^{+}$ do
4       remove segment $\theta_{i}$ from $\mathcal{J}$
5let $\mathcal{J}^{\prime}=((\beta_{0},\sigma_{0}),\dots,(\beta_{t},\sigma_{t}))$ be the preschedule after these modification steps. for $i=0,\dots,t$ do let $\hat{\mu}_{\beta_{i}}\leftarrow{\tt Sample}(\beta_{i};\tfrac{\nu}{2},e^{-\nu/2}w_{i},\frac{\gamma}{4(t+1)})$ where $w_{i}=\phi/|\sigma_{i}|$
6 for $i=1,\dots,t$ do
7       if  $\exists\;k\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$ with  $\hat{\mu}_{\beta_{i-1}}(k)\geq e^{-\nu/2}w_{i-1}$ and $\hat{\mu}_{\beta_{i}}(k)\geq e^{-\nu/2}w_{i}$  then
8             set $k_{i}=k$ for arbitrary such $k$
9      else return ERROR.
return covering schedule $\mathcal{I}=(\beta_{0},e^{-\nu}w_{0},k_{1},\beta_{1},e^{-\nu}w_{1},k_{2},\dots,k_{t},\beta_{t},e^{-\nu}w_{t})$
Algorithm 12 ${\tt TryToUncross}(\cal J,\gamma)$ for preschedule $\mathcal{J}$.
Proposition 38.
At Line 5, the preschedule $\mathcal{J}^{\prime}$ satisfies $\sum_{i=0}^{t}\frac{1}{w_{i}}\leq\frac{2(n+1)}{\phi}$. Furthermore, if the original preschedule $\mathcal{J}$ is $\phi$-proper, then so is $\mathcal{J}^{\prime}$.
Proof.
When segments are merged at Line 2, the new segment has a larger span, so it preserves properness. The discarding step at Line 4 clearly does not make a schedule improper.
We claim that for any $k\in\mathcal{H}$, there are at most two segments $\theta_{i}=(\beta_{i},\sigma_{i})\in\mathcal{J}^{\prime}$ with $k\in\sigma_{i}$. For, suppose $k\in\sigma_{i_{1}}\cap\sigma_{i_{2}}\cap\sigma_{i_{3}}$ with $i_{1}<i_{2}<i_{3}$. Then by (I2), we have $\sigma_{i_{2}+1}^{-}\leq\sigma_{i_{3}}^{-}\leq k$ and $\sigma_{i_{2}-1}^{+}\geq\sigma_{i_{1}}^{+}\geq k$. So $\sigma_{i_{2}+1}^{-}\leq\sigma_{i_{2}-1}^{+}$ and we could have discarded segment $i_{2}$ in the loop at Line 3.
Consequently, we calculate $\sum_{i=0}^{t}\frac{1}{w_{i}}=\frac{1}{\phi}\sum_{i=0}^{t}|\sigma_{i}|=\frac{1}{\phi}\sum_{k\in\mathcal{H}}|\{i:k\in\sigma_{i}\}|\leq\frac{2(n+1)}{\phi}.$
Theorem 39.
The algorithm TryToUncross satisfies the following properties:
(a) The output is either ERROR or a covering schedule $\mathcal{I}$ with ${\tt InvWt}(\mathcal{I})\leq\frac{2e^{\nu}(n+1)}{\phi}$.
(b) The output is an improper covering schedule with probability at most $\gamma$.
(c) If $\mathcal{J}$ is $\phi$-proper, then it outputs ERROR with probability at most $\gamma$.
(d) The cost is $O(n\rho\log\frac{n}{\gamma})$.
Proof.
From Proposition 38, we have ${\tt InvWt}(\mathcal{I})=\sum_{i=0}^{t}\frac{1}{e^{-\nu}w_{i}}\leq\frac{2e^{\nu}(n+1)}{\phi}$. Similarly, the algorithm cost from Line 6 is $\sum_{i=0}^{t}O(w_{i}\log\tfrac{t}{\gamma})\leq O(\frac{2(n+1)}{\phi}\cdot\log\tfrac{n}{\gamma})\leq O(n\rho\log\tfrac{n}{\gamma})$, thus showing part (d).
To show $\mathcal{I}$ is a covering schedule, we need to show that $k_{1}<\dots<k_{t}$ strictly and $\beta_{0}=\beta_{\min}<\dots<\beta_{t}=\beta_{\max}$ strictly. The condition $\beta_{0}=\beta_{\min},\beta_{t}=\beta_{\max}$ follows from (I1) while the bound $\beta_{i}<\beta_{i+1}$ follows from the loop at Line 1. To show $k_{i}<k_{i+1}$ for $i\in\{1,\dots,t-1\}$, observe that from (I0) we have $\sigma_{i-1}^{+}\geq\sigma_{i}^{-}$ and $\sigma_{i}^{+}\geq\sigma_{i+1}^{-}$. Due to the loop at Line 3, we must have $k_{i+1}\geq\sigma_{i+1}^{-}>\sigma_{i-1}^{+}\geq k_{i}$ (else segment $\theta_{i}$ would have been discarded).
Now suppose that Line 6 well-estimates each $\sigma_{i}^{+},\sigma_{i}^{-},\sigma_{i+1}^{-},\sigma_{i+1}^{+}$; by specification of the parameters this has probability at least $1-\gamma$. We claim that the algorithm outputs either a proper covering schedule or ERROR, and that the latter case only holds if $\mathcal{J}^{\prime}$, and hence $\mathcal{J}$, is improper.
First, if the algorithm reaches Line 11, then $\hat{\mu}_{\beta_{i}}(k_{i})\geq e^{-\nu/2}w_{i}$ and $\hat{\mu}_{\beta_{i-1}}(k_{i})\geq e^{-\nu/2}w_{i-1}$ for each $i$. Since Line 6 well-estimates $k_{i}$, this implies $\mu_{\beta_{i}}(k_{i})\geq e^{-\nu}w_{i}$ and $\mu_{\beta_{i-1}}(k_{i})\geq e^{-\nu}w_{i-1}$. So $\mathcal{I}$ is proper.
Next, suppose $\mathcal{J}^{\prime}$ is $\phi$-proper but the algorithm outputs ERROR at an iteration $i$. By definition of $\phi$-properness, we have $\mu_{\beta_{i-1}}(\sigma_{i-1}^{+})\geq w_{i-1}$ and $\mu_{\beta_{i}}(\sigma_{i}^{-})\geq w_{i}$. Since Line 6 well-estimates $\sigma_{i-1}^{+}$ and $\sigma_{i}^{+}$, this implies $\hat{\mu}_{\beta_{i-1}}(\sigma_{i-1}^{+})\geq e^{-\nu/2}w_{i-1}$ and $\hat{\mu}_{\beta_{i}}(\sigma_{i}^{-})\geq e^{-\nu/2}w_{i}$. Neither value $k\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$ satisfied the check at Line 8, so $\hat{\mu}_{\beta_{i}}(\sigma_{i-1}^{+})<e^{-\nu/2}w_{i}$ and $\hat{\mu}_{\beta_{i-1}}(\sigma_{i}^{-})<e^{-\nu/2}w_{i-1}$. In turn, since Line 6 well-estimates these values, we have $\mu_{\beta_{i}}(\sigma_{i-1}^{+})<w_{i}$ and $\mu_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}$. But now $\mu_{\beta_{i}}(\sigma_{i-1}^{+})\mu_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}w_{i}\leq\mu_{\beta_{i}}(\sigma_{i}^{-})\mu_{\beta_{i-1}}(\sigma_{i-1}^{+})$. This contradicts the Uncrossing Inequality since $\sigma_{i}^{-}\leq\sigma_{i-1}^{+}$ by (I0). ∎
We can finish by combining all the preschedule processing algorithms, as follows:
1 while true do
2       call Algorithm 11 with appropriate constants $\nu,\lambda,\tau$ to compute preschedule ${\cal J}$
3       call ${\cal I}\leftarrow{\tt TryToUncross}({\cal J},\gamma/4)$
4       if $\mathcal{I}\neq$ ERROR then return $\mathcal{I}$
5
Algorithm 13 Algorithm ${\tt FindCoveringSchedule}(\gamma)$
By Lemma 32 and Theorem 39, each iteration of Algorithm 13 terminates with probability at least $\tfrac{1}{2}(1-\gamma/4)\geq 3/8$, so there are $O(1)$ expected iterations. Each call to TryToUncross has cost $O(n\rho\log\frac{n}{\gamma})$. By Proposition 31, each call to Algorithm 11 has cost $O(n\log q+n\rho\log^{2}n)$.
By Theorem 39(a), we have ${\tt InvWt}(\mathcal{I})\leq 2\rho(n+1)\cdot e^{\nu}\lambda^{3}/\tau$. The term $e^{\nu}\lambda^{3}/\tau$ gets arbitrarily close to $2$ for constants $\nu,\lambda,\tau$ sufficiently close to $0,1,\frac{1}{2}$ respectively.
Finally, by Proposition 39, each iteration of Algorithm 13 returns a non-proper covering schedule with probability at most $\gamma/4$ (irrespective of the choice of $\mathcal{J}$). Thus, the total probability of returning a non-proper covering schedule over all iterations is at most $\sum_{i=0}^{\infty}(3/8)^{i}\gamma/4=2\gamma/5$.
This shows Theorem 30.
7 Combinatorial applications
[…]
We will examine three combinatorial problems from graph theory: connected subgraphs, matchings, and independent sets. There is essentially only one piece of information we need about the underlying Gibbs distribution, namely, the value $\Delta_{\min}$. Esentially all other problem-specific details can be ignored. Furthermore, for log-concave distributions, there is a simple “automatic” bound on $\Delta_{\min}$; combined with our algorithm for log-concave distributions, this gives a particularly clean algorithmic result.
Theorem 40.
Suppose the counts are non-zero and log-concave. If $\beta_{\min}\leq\log\frac{c_{0}}{c_{1}}$ and $\beta_{\max}\geq\log\frac{c_{n-1}}{c_{n}}$, then $\Delta_{\min}\geq\frac{1}{n+1}$ and $\log Q(\beta_{\max})\leq q=3n\Gamma$ for $\Gamma=\max\{\beta_{\max},\log\frac{c_{1}}{c_{0}},1\}$.
In particular, for $\delta=\frac{1}{n+1}$, we can solve ${P^{\delta,\varepsilon}_{\tt count}}$ with cost $O\bigl{(}\min\bigl{\{}\frac{n\Gamma\log n\log\frac{1}{\gamma}}{\varepsilon^{2}},\frac{n^{2}\log\frac{1}{\gamma}}{\varepsilon^{2}}+n\log\Gamma\bigr{\}}\bigr{)}$.
Proof.
Define $b_{i}=\log(c_{i-1}/c_{i})$ for $i=1,\dots,n$; the sequence $b_{1},\dots,b_{n}$ is non-decreasing since $c_{i}$ is log-concave. We claim that for each $i,k\in\mathcal{H}$, there holds
$c_{i}e^{ib_{i}}\geq c_{k}e^{kb_{i}}$ (22)
To show this for $k>i$, we note that $\frac{c_{i}e^{ib_{i}}}{c_{k}e^{kb_{i}}}=e^{(i-k)b_{i}}\prod_{j=i}^{k-1}\frac{c_{j}}{c_{j+1}}=e^{\sum_{j=i}^{k-1}b_{j+1}-b_{i}}\geq 1$. A similar calculation holds for $k<i$. Since $\mu_{\beta}(k)\propto c_{k}e^{\beta k}$, Eq. (22) implies
$\mu_{b_{i}}(i)=\frac{\mu_{b_{i}}(i)}{\sum_{k=0}^{n}\mu_{b_{i}}(k)}\geq\frac{\mu_{b_{i}}(i)}{\sum_{k=0}^{n}\mu_{b_{i}}(i)}=\frac{1}{n+1}.$
Similarly, we have $\mu_{b_{1}}(0)\geq\frac{1}{n+1}$. So $b_{i}\in[b_{1},b_{n}]\subseteq[\beta_{\min},\beta_{\max}]$ and $\Delta_{\min}\geq\tfrac{1}{n+1}$ as claimed. We next turn to the bound on $Q(\beta_{\max})$. We have the lower bound $Z(\beta_{\min})\geq c_{0}$. To upper-bound $Z(\beta_{\max})$, we observe from Eq. (22) and from our hypothesis $\beta_{\max}\geq b_{n}$ that for every $k\leq n$, we have
$\frac{c_{n}e^{n\beta_{\max}}}{c_{k}e^{k\beta_{\max}}}=\frac{c_{n}e^{nb_{n}}}{c_{k}e^{kb_{n}}}\cdot e^{(\beta_{\max}-b_{n})(n-k)}\geq 1.$
So $Z(\beta_{\max})=\sum_{i}c_{i}e^{i\beta_{\max}}\leq(n+1)c_{n}e^{n\beta_{\max}}$. This implies $Q(\beta_{\max})\leq\frac{e^{n\beta_{\max}}(n+1)c_{n}}{c_{0}}$. By telescoping products, $\frac{c_{n}}{c_{0}}=\prod_{i=1}^{n}\frac{c_{i}}{c_{i-1}}\leq(\frac{c_{1}}{c_{0}})^{n}$, giving $Q(\beta_{\max})\leq e^{n\beta_{\max}}\cdot(n+1)\cdot(\frac{c_{1}}{c_{0}})^{n}\leq e^{n\Gamma}\cdot(n+1)\cdot e^{n\Gamma}\leq e^{q}$. ∎
7.1 Counting connected subgraphs
Let $G=(V,E)$ be a connected graph. The problem of network reliability is to sample a connected subgraph $G^{\prime}=(V,E^{\prime})$ with probability proportional to $\prod_{f\in E^{\prime}}(1-p(f))\prod_{f\in E-E^{\prime}}p(f)$, for any weighting function $p:E\rightarrow[0,1]$. This can be interpreted as each edge $f$ “failing” independently with probability $p(f)$, and conditioning on the resulting subgraph remaining connected. Equivalently, if we set $p(f)=\frac{e^{\beta}}{1+e^{\beta}}$ for all edges $f$, it can be interpreted as a Gibbs distribution where $c_{i}$ is the number of connected subgraphs of $G$ with $|E|-i$ edges and with $n=|E|-|V|+1$.
Note that $c_{i}$ is the number of independent sets in the co-graphic matroid. By the result of [1], this sequence $c_{i}$ is log-concave. Anari et al. [2] provides an FPRAS for counting independent sets in arbitrary matroids, which would include connected subgraphs, but do not provide concrete complexity estimates for their algorithm. There have also been a series of exact and approximate sampling algorithms for network reliability via Gibbs distributions [12, 11, 7]. The most recent result is due to Chen, Zhang, & Zou [7] which we summarize as follows:
Theorem 41 ([7], Corollary 3).
There is an algorithm to approximately sample from the network reliability distribution for any value of $\beta$, up to total variation distance $\rho$, with expected runtime $O(|E|(e^{\beta}+1)\log(|V|/\rho)\log^{2}|V|)$.
Based on Theorem 41, we can immediately get our counting algorithm.
Proof of Theorem 7.
Here $c_{0}=1$ and $c_{1}/c_{0}\leq|E|$ and $c_{n-1}/c_{n}\leq|E|$. So we can apply Theorem 40, setting $\beta_{\max}=\log|E|,\beta_{\min}=-\infty$ and $\Gamma=\log|E|$. The algorithm uses $O(\frac{|E|\log^{2}|E|}{\varepsilon^{2}})$ samples in expectation to achieve failure probability $\gamma=O(1)$. Accordingly, we need to run the approximate sampler of Theorem 41 with $\rho=\operatorname*{poly}(n,1/\varepsilon)$. The total algorithm runtime is then $\tilde{O}(|E|^{3}/\varepsilon^{2})$. ∎
7.2 Counting matchings in high-degree graphs
Consider a graph $G=(V,E)$ with a perfect matching. For $i=0,\dots,|V|/2$, let $c_{i}$ denote the number of $i$-edge matchings. As originally shown in [13], the sequence $c_{i}$ is log-concave. In [18, 19], Jerrum & Sinclair described an MCMC algorithm to approximately sample from the Gibbs distribution on matchings. We quote their most optimized result as follows:
Theorem 42 ([19]).
There is an algorithm to approximately sample from the Gibbs distribution at $\beta$ with up to total variation distance $\rho$, with expected runtime $\tilde{O}(|E||V|^{2}(1+e^{\beta})\log\frac{1}{\rho})$.
Proof of Theorem 8.
Here $c_{0}=1$ and $n=|V|/2$ and $c_{1}/c_{0}\leq|E|$, and by assumption $c_{n-1}/c_{n}\leq f$. So we can apply Theorem 40 setting $\beta_{\min}=-\log|E|,\beta_{\max}=\log f$, and $\Gamma\leq\max\{\log|E|,\log f\}$. The algorithm uses $O(n\log(|E|f)\log^{2}n/\varepsilon^{2})$ samples in expectation for failure probability $\gamma=O(1)$.
By Theorem 55, we can take $\rho=\operatorname*{poly}(1/n,f,\varepsilon)$ to ensure that the sampling oracle is sufficiently close to the Gibbs distribution. Then Theorem 42 requires $O(|E||V|^{2}f\operatorname*{polylog}(|V|,f,1/\varepsilon))$ time per sample, and overall our FPRAS has runtime of $\tilde{O}(\frac{|E||V|^{3}f}{\varepsilon^{2}})$.
Note also that, as shown in [18], if $G$ has minimum degree at least $|V|/2$ then $c_{|V|/2}>0$ and $c_{|V|/2-1}/c_{|V|/2}\leq O(|V|^{2})$ and clearly $|E|\leq O(|V|^{2})$. ∎
7.3 Counting independent sets in bounded-degree graphs
Let $G=(V,E)$ be a graph of of maximum degree $D$. A key problem in statistical physics is to efficiently sample the independent sets of $G$. We can consider the Gibbs distribution on independent sets weighted by their size, i.e. each independent set $I\subseteq V$ is sampled with probability $\lambda^{|I|}=e^{\beta|I|}$ for a given parameter $\lambda=e^{\beta}$, and $c_{k}$ is the number of independent sets of cardinality $k$. Note that the Gibbs distribution is not necessarily log-concave.
As shown in [8], this distribution has a critical hardness threshold defined by $\lambda_{c}=\frac{(D-1)^{D-1}}{(D-2)^{D}}\approx e/D$. It is intractable to sample from the Gibbs distribution beyond $\beta>\log\lambda_{c}$ and there is a polynomial-time sampler for the Gibbs distribution for $\beta<\log\lambda_{c}$. We quote the following result:
Theorem 43 ([8]).
Let $D\geq 3$ and $\xi>0$ be any fixed constants. There is an algorithm to approximately sample from the Gibbs distribution at $\beta\in[-\infty,\log(\lambda_{c})-\xi]$, up to total variation distance $\rho$, with runtime $O(|V|\log|V|\log\tfrac{1}{\rho})$. [^4] Concretely, [8] describes an MCMC known as Glauber dynamics whose stationary distribution is the Gibbs distribution. They show this Markov chain has mixing time $O(|V|\log|V|)$. For a constant-degree graph, each step of the Glauber dynamics can be executed in $O(1)$ time.
The related problem of estimating counts was considered in [9, 17]. They identified a related threshold value $\alpha_{c}=\frac{\lambda_{c}}{1+(D+1)\lambda_{c}}$ such that for $k>\alpha_{c}|V|$, it is intractable to estimate $c_{k}$ or sample approximately uniformly from size-$k$ independent sets, while on the other hand, for any constants $D\geq 3,\xi>0$ and any $k<(\alpha_{c}-\xi)|V|$, there is a polynomial-time algorithm for both tasks. A key technical result shown in [17] was that, in the latter regime, $\mu_{\beta}$ could be approximated by a normal distribution, i.e. it obeyed a type of Central Limit Theorem.
For our purposes, using Theorem 3.1 of [17], we can get a cruder estimate:
Lemma 44 ([17]).
Let $D\geq 3$ and $\xi>0$ be any fixed constants. There is a constant $\xi^{\prime}>0$ such that, for any $k\leq(\alpha_{c}-\xi)|V|$, there is a value $\beta<\log(\lambda_{c})-\xi^{\prime}$ with $\mu_{\beta}(k)\geq\Omega(1/\sqrt{|V|})$.
This bound alone is all we need for our counting algorithm, which we summarize as follows:
Proof of Theorem 9.
Here $c_{0}=1$ and $n=|V|$. We set $\beta_{\min}=-\infty$ and $\beta_{\max}=\log(\lambda_{c})-\xi^{\prime}$, so $\frac{Q(\beta_{\max})}{Q(\beta_{\min})}\leq\frac{2^{n}e^{\beta_{\max}n}}{c_{0}}$; in particular, since $\beta_{\max}=O(1)$ (for fixed $D$), we can take $q=\Theta(n)$. It suffices to solve ${P^{\delta,\varepsilon}_{\tt count}}$ for $\delta=\Omega(1/\sqrt{n})$. We use the continuous-setting algorithm of Theorem 22; for $\gamma=O(1)$, it has expected sample complexity $O\bigl{(}\frac{q\log n+(\sqrt{q\log n}/\delta)\log\tfrac{q}{\delta}}{\varepsilon^{2}}\bigr{)}=O\bigl{(}\frac{n\log^{3/2}n}{\varepsilon^{2}}\bigr{)}$. The sampler of Theorem 43 then should set $\rho=\operatorname*{poly}(n,1/\varepsilon)$, and the overall runtime is $O\bigl{(}\frac{n^{2}\log^{5/2}n\log(n/\varepsilon)}{\varepsilon^{2}}\bigr{)}=\tilde{O}(n^{2}/\varepsilon^{2}).$
8 Lower bounds on sample complexity
In this section, we show lower bounds for the sample complexity needed to solve ${P^{{\varepsilon}}_{\tt ratio-pt}}$ and ${P^{\delta,\varepsilon}_{\tt count}}$ in the black-box model. The specification of problem ${P^{\delta,\varepsilon}_{\tt count}}$, while useful from an algorithmic point of view, has some restrictive choices for its normalization and error bounds; to get more general lower bounds for count estimation, we consider a relaxed problem called ${P^{\delta,\varepsilon}_{\tt count-weak}}$, namely, to compute a vector $\hat{c}\in\mathbb{R}_{\geq 0}^{\mathcal{F}}$ satisfying the following property:
$|\log\frac{\hat{c}_{x}}{\hat{c}_{y}}-\log\frac{c_{x}}{c_{y}}|\leq\varepsilon\qquad\qquad\text{ for all pairs $x,y$ with $\Delta(x),\Delta(y)\geq\delta$}$
One can easily check that a solution $\hat{\pi}$ to ${P^{\delta,\varepsilon/5}_{\tt count}}$ is also a solution to ${P^{\delta,\varepsilon}_{\tt count-weak}}$. Thus, up to constant factors in parameters, the problem ${P_{\tt count-weak}}$ is easier to solve compared to ${P_{\tt count}}$.
Our strategy for the lower bounds is to construct a target instance $c$ surrounded by an envelope of alternate instances, such that solving ${P^{{\varepsilon}}_{\tt ratio-pt}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$ on an unknown instance $\tilde{c}$ allows us to distinguish between whether $\tilde{c}$ is an alternate instance or the base instance. Specifically, let us choose $d$ vectors $\lambda^{(1)},\dots,\lambda^{(d)}\in\mathbb{R}^{\mathcal{F}}$. We will define $2d$ alternate instances $r=-d,\dots,-1,1,\dots,d$ which straddle the base instance $c=c^{(0)}$, defined as:
[…]
For $r=-d,\dots,d$, we define $\mu_{\beta}^{(r)}$ to be the Gibbs distribution for instance $c^{(r)}$, and $Z^{(r)}$ to be its partition function, and $z^{(r)}=\log Z^{(r)}$, etc. When the superscript is omitted, we always refer to the base instance $c=c^{(0)}$, e.g. we write $z(\beta)=z^{(0)}(\beta),\Delta(x)=\max_{\beta\in[\beta_{\min},\beta_{\max}]}\mu_{\beta}^{(0)}(x)$.
We also define parameters
$\kappa=\max_{\begin{subarray}{c}x\in\mathcal{F},r\in\{1,\dots,d\}\end{subarray}}|\lambda^{(r)}_{x}|,\qquad\Psi=\max_{\beta\in[\beta_{\min},\beta_{\max}]}\sum_{r=1}^{d}\log\frac{Z^{(+r)}(\beta)Z^{(-r)}(\beta)}{Z(\beta)^{2}}.$
In this setting, we have a key lower bound on the sample complexity of any procedure to distinguish the distributions.
Lemma 45 (Indistinguishability lemma).
Let $\mathfrak{A}$ be an algorithm which generates queries $\beta_{1},\dots,\beta_{T}\in[\beta_{\min},\beta_{\max}]$ and receives values $x_{1},\dots,x_{T}$, where each $x_{i}$ is drawn from $\mu_{\beta_{i}}$. At some point the procedure stops and outputs TRUE or FALSE. The queries $\beta_{i}$ and the stopping time $T$ may be adaptive and may be randomized.
Let $\gamma\in(0,1/4)$. If $\mathfrak{A}$ outputs FALSE on input $c^{(0)}$ with probability at least $1-\gamma$ and outputs FALSE on inputs $c^{(-d)},\dots c^{(-1)},c^{(1)},\dots,c^{(d)}$ with probability at least $1-\gamma$, then $\mathfrak{A}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ when run on instance $c^{(0)}$.
Proof.
Any finite run of $\mathfrak{A}$ can be described by a random vector $X=((\beta_{1},x_{1}),\ldots,(\beta_{T},x_{T}),\sigma)$, where $\sigma\in\{\text{TRUE},\text{FALSE}\}$. The probability measure ${\mathbb{P}}^{(i)}(\cdot)$ over such runs $X$ on an instance $c^{(r)}$ can be decomposed as:
$\operatorname{d}\!{{\mathbb{P}}}^{(r)}(X)=\psi^{(r)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X$ (23)
where the measure $\mu^{\mathfrak{A}}$ depends only on the algorithm $\mathfrak{A}$, and the function $\psi^{(r)}$ is defined via
$\psi^{(r)}((\beta_{1},x_{1}),\ldots,(\beta_{T},x_{T}),\sigma)=\prod_{i=1}^{T}\mu_{\beta_{i}}^{(r)}(x_{i})$
Define $R=\{-d,\dots,-1,+1,\dots,+d\}$ (the alternate problem instances). For any finite run $X=((\beta_{1},x_{1}),\ldots,(\beta_{T},x_{T}),\sigma)$ we have
$\prod_{r\in R}\psi^{(r)}(X)$ $=\prod_{r\in R}\prod_{t=1}^{T}\mu^{(r)}_{\beta_{t}}(x_{t})=\prod_{t=1}^{T}(\mu_{\beta_{t}}(x_{t}))^{2d}\prod_{r\in R}\frac{c^{(r)}_{x_{t}}e^{\beta_{t}x_{t}}Z(\beta_{t})}{c_{x_{t}}e^{\beta_{t}x_{t}}Z^{(r)}(\beta_{t})}$
$=(\psi^{(0)}(X))^{2d}\prod_{t=1}^{T}\prod_{r=1}^{d}\frac{Z(\beta_{t})^{2}}{Z^{(+r)}(\beta_{t})Z^{(-r)}(\beta_{t})}\geq(\psi^{(0)}(X))^{2d}e^{-T\Psi}$ (24)
Let $\mathcal{X}$ denote the set of runs of length at most $\tau=\frac{2d\log(1/(2\gamma))}{\Psi}$. Partition $\mathcal{X}$ into sets of runs $\mathcal{X}^{\text{T}},\mathcal{X}^{\text{F}}$ which output TRUE and FALSE respectively. For any run $X\in\mathcal{X}$ of length $T$, Eq. (24) gives:
$\frac{1}{2d}\sum_{r\in R}\psi^{(r)}(X)\geq\Bigl{(}\prod_{r\in R}\psi^{(r)}(X)\Bigr{)}^{1/(2d)}\geq\psi^{(0)}(X)e^{-T\Psi/2d}\geq 2\gamma\psi^{(0)}(X)$ (25)
where the first bound comes from the inequality between arithmetic and geometric means, and the last inequality holds since $T\leq\tau$.
By hypothesis of the lemma, we have ${\mathbb{P}}^{(r)}(\mathcal{X}^{\text{T}})<\gamma$ for all $r\in R$. Using Eq. (25), this implies
${\mathbb{P}}^{(0)}(\mathcal{X}^{\text{T}})=\negthinspace\int_{{\cal X}^{\text{T}}}\negthinspace\psi^{(0)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X\leq\frac{1}{4d\gamma}\sum_{r\in R}\int_{{\cal X}^{\text{T}}}\negthinspace\psi^{(r)}(X)\>\operatorname{d}\!{\mu}^{\mathfrak{A}}X=\frac{1}{4d\gamma}\sum_{r\in R}{\mathbb{P}}^{(r)}(\mathcal{X}^{\text{T}})\leq\frac{1}{4d\gamma}\sum_{r\in R}\gamma=\frac{1}{2}.$
By hypothesis, we have ${\mathbb{P}}^{(0)}(\mathcal{X}^{\text{F}})<\gamma$. So ${\mathbb{P}}^{(0)}(T>\tau)=1-{\mathbb{P}}^{(0)}(\mathcal{X}^{\text{T}})-{\mathbb{P}}^{(0)}(\mathcal{X}^{\text{F}})\geq 1-\gamma-\frac{1}{2}\geq\frac{1}{4}$. The expected cost of $\mathfrak{A}$ on $c^{(0)}$ is at least $\frac{1}{4}\cdot\tau=\Omega(\frac{d\log(1/\gamma)}{\Psi})$. ∎
Corollary 46.
Suppose that $z^{(0)}(\beta_{\min},\beta_{\max})\leq q-2\kappa$.
(a) If $|z(\beta_{\min},\beta_{\max})-z^{(r)}(\beta_{\min},\beta_{\max})|>2\varepsilon$ for all $r=\pm 1,\dots,\pm d$, then any algorithm for ${P^{{\varepsilon}}_{\tt ratio-pt}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ on $c$.
(b) If for each $r=1,\dots,d$ there are $x,y\in\mathcal{F}$ with $\Delta(x),\Delta(y)\geq\delta e^{2\kappa}$ and $|\lambda^{(r)}_{x}-\lambda^{(r)}_{y}|>2\varepsilon$, then any algorithm for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})$ on $c$. (We refer to $x,y$ as the witnesses for index $r$).
Proof.
As $|\log c^{(r)}_{x}-\log c_{x}|\leq\kappa$ for all $x$, we have $\frac{Z^{(r)}(\beta)}{Z(\beta)}\in[e^{-\kappa},e^{\kappa}]$ and $\mu^{(r)}_{\beta}(x)=\frac{c_{x}^{(r)}e^{\beta x}}{Z^{(r)}(\beta)}\geq\mu_{\beta}(x)e^{-2\kappa}$. So $z^{(r)}(\beta_{\min},\beta_{\max})\leq q$ for all $r=\pm 1,\dots,\pm d$. In particular, algorithms for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ or ${P^{{\varepsilon}}_{\tt ratio-pt}}$, which take as input the upper-bound value $q$, must also work for all instances $c^{(-d)},\dots,c^{(+d)}$. Furthermore, such algorithms can be used to distinguish $c^{(0)}$ from $c^{(\pm 1)},\dots,c^{(\pm d)}$. Namely, given a solution $\hat{Q}(\beta_{\max})$ for ${P^{{\varepsilon}}_{\tt ratio-pt}}$, we output TRUE if and only if $|\log\hat{Q}(\beta_{\max})-\log Q^{(0)}(\beta_{\max})|\leq\varepsilon$. Likewise, given a solution $\hat{c}$ to ${P^{\delta,\varepsilon}_{\tt count-weak}}$, we output TRUE if and only if for all $r=1,\dots,d$ the witnesses $x,y$ satisfy $|\log\frac{\hat{c}_{x}}{\hat{c}_{y}}-\log\frac{c_{x}}{c_{y}}|\leq\varepsilon$. Note that $\Delta^{(\pm r)}(x),\Delta^{(\pm r)}(y)\geq\delta$ and so $x,y$ can serve as witnesses for both instances $c^{(+r)}$ and $c^{(-r)}$. ∎
By applying Corollary 46 to carefully constructed instances, we will show the following results:
Theorem 47.
Let $n\geq n_{0},q\geq q_{0},\varepsilon<\varepsilon_{0},\delta<\delta_{0},\gamma<1/4$ for certain absolute constants $n_{0},q_{0},\varepsilon_{0},\delta_{0}$. There are problem instances $\mu^{(0)},\dots\mu^{(d)}$ which satisfy the given bounds $n$ and $q$ such that:
(a) ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{\min\{q+\sqrt{q}/\delta,n^{2}+n/\delta\}}{\varepsilon^{2}}\log\frac{1}{\gamma})$, and $\mu$ is integer-valued.
(b) ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{1/\delta+\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma})$, and $\mu$ is log-concave.
(c) ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega(\frac{\min\{q,n^{2}\}}{\varepsilon^{2}}\log\frac{1}{\gamma})$, and $\mu$ is log-concave.
(d) ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{(q+\sqrt{q}/\delta)}{\varepsilon^{2}}\log\frac{1}{\gamma})$.
(e) ${P^{{\varepsilon}}_{\tt ratio-pt}}$ requires cost $\Omega(\frac{q}{\varepsilon^{2}}\log\frac{1}{\gamma})$.
Theorem 47 immediately implies Theorems 2 and 4.
The lower bound in terms of $\delta$ for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ is rather trivial, and does not really depend on properties of Gibbs distributions. We summarize it in the following result:
Proposition 48.
For $\delta<1/10$ and $\varepsilon<1/20$, there is a log-concave instance with arbitrary $n\geq 1$ and $\beta_{\min}=\beta_{\max}=0$, for which solving ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{\log(1/\gamma)}{\delta\varepsilon^{2}})$.
Proof.
Take $d=1$ with $c_{0}=2\delta,c_{1}=1,c_{2}=\dots=c_{n}=0$, and $\lambda^{(1)}_{0}=3\varepsilon,\lambda^{(1)}_{1}=\dots=\lambda^{(1)}_{n}=0$. We can apply Corollary 46(b) with witnesses $x=0,y=1$. It is straightforward to calculate $\Delta(0),\Delta(1)\geq 5/3\cdot\delta\geq\delta e^{2\kappa}$ and $\Psi=\log\bigl{(}\frac{(2\delta e^{-3\varepsilon}+1)(2\delta e^{3\varepsilon}+1)}{(2\delta+1)^{2}}\bigr{)}\leq O(\delta\varepsilon^{2})$. ∎
We next turn to the non-trivial lower bounds in terms of $n$ and $q$.
8.1 Bounds in the log-concave setting
For this construction, take $d=1,\beta_{\min}=0,\beta_{\max}=n$, and define $c_{0},\dots,c_{n}$ to be the coefficients of the polynomial $g(x)=\prod_{k=0}^{n-1}(e^{k}+x)$, and set $\lambda^{(1)}_{k}=k\nu$ for $k=0,\dots,n$ where $\nu=10\varepsilon/n$. Since $g(x)$ is a real-rooted polynomial, and counts $c^{(\pm 1)}$ are derived by an exponential shift, the counts $c^{(0)},c^{(+1)},c^{(-1)}$ are log-concave [5].
Here $\prod_{k=0}^{n-1}(e^{k}+x)=c_{0}+c_{1}x+\dots+c_{n}x^{n}$, and $c_{0}=\prod_{k=0}^{n-1}e^{k}$ and $c_{n}=1$, and we have
$Z(\beta)=\sum_{x}c_{x}e^{\beta x}=g(e^{\beta x})=\prod_{k=0}^{n-1}(e^{k}+e^{\beta}),\quad\text{and}\quad Z^{(\pm 1)}(\beta)=\sum_{x}c_{x}e^{\pm\nu}e^{\beta x}=Z(\beta\pm\nu)$
For intuition, consider independent random variables $X_{0},\dots,X_{n-1}$, wherein each $X_{i}$ is Bernoulli-$p_{i}$ for $p_{i}=\frac{e^{\beta}}{e^{i}+e^{\beta}}$. Then $\mu_{\beta}$ is the probability distribution on random variable $X=X_{0}+\dots+X_{n-1}$. Distributions $\mu^{(\pm 1)}$ are derived by slightly shifting the probabilities $p_{i}$.
Observation 49.
For any values $\beta,x$ we have $|z(\beta+x)-z(\beta)-z^{\prime}(\beta)x|\leq 2x^{2}$.
Proof.
We bound the second derivative of function $z$ as
$z^{\prime\prime}(\beta)=\sum_{k=0}^{n-1}\frac{e^{k}e^{\beta}}{(e^{k}+e^{\beta})^{2}}\leq\sum_{k\leq\beta}\frac{e^{k}e^{\beta}}{e^{2\beta}}+\sum_{k\geq\beta}\frac{e^{k}e^{\beta}}{e^{2k}}\leq\frac{e}{e-1}+\frac{e}{e-1}\leq 4.$
The result then follows from Taylor’s theorem. ∎
Lemma 50.
For $n\geq 7$ and $\varepsilon<1/100$, the following bounds hold:
(a) $z(\beta_{\min},\beta_{\max})\leq n^{2}$.
(b) $|z^{(r)}(\beta_{\min},\beta_{\max})-z(\beta_{\min},\beta_{\max})|>2\varepsilon$ for $r=\pm 1$.
(c) $\Psi\leq O(\varepsilon^{2}/n^{2}).$
(d) $\mu_{0}(0)\geq 1/5$ and $\mu_{\beta_{\max}}(n)\geq 1/5$.
Proof.
For (a), we calculate:
$z(0,n)=\sum_{k=0}^{n-1}\log(e^{k}+e^{n})-\sum_{k=0}^{n-1}\log(e^{k}+1)=\sum_{k=0}^{n-1}\log\bigl{(}\frac{e^{k}+e^{n}}{e^{k}+1}\bigr{)}\leq\sum_{k=0}^{n-1}\log\bigl{(}\frac{e^{n}}{1}\bigr{)}=n^{2}.$
For (b), we can use Observation 49 and the formula $z^{(\pm 1)}(\beta)=z(\beta\pm\nu)$ to get:
$|z^{(\pm 1)}(\beta_{\min},\beta_{\max})-z(\beta_{\min},\beta_{\max})|=\bigl{|}\bigl{(}z(n\pm\nu)-z(n)\bigr{)}-\bigl{(}z(\pm\nu)-z(0)\bigr{)}\bigr{|}$
$\qquad\qquad\geq(z^{\prime}(n)\nu-2\nu^{2})-(z^{\prime}(0)\nu+2\nu^{2})=(z^{\prime}(n)-z^{\prime}(0))\nu-4\nu^{2}$
We have $z^{\prime}(n)=\sum_{k=0}^{n-1}\frac{e^{n}}{e^{k}+e^{n}}\geq\sum_{k=0}^{n-1}\frac{e^{n}}{e^{n}+e^{n}}=n/2$ and $z^{\prime}(0)=\sum_{k=0}^{n-1}\frac{1}{1+e^{k}}\leq\sum_{k=0}^{\infty}e^{-k}=\frac{e}{e-1}$. So $|z^{(+\pm)}(\beta_{\min},\beta_{\max})-z(\beta_{\min},\beta_{\max})|\geq(n/2-\frac{e}{e-1})\nu-4\nu^{2}$, which is larger than $2\varepsilon$ for $n\geq 7,\varepsilon<1/100$.
For (c), we have $\log\frac{Z^{(+1)}(\beta)Z^{(-1)}(\beta)}{Z(\beta)^{2}}=(z(\beta+\nu)-z(\beta))+(z(\beta-\nu)-z(\beta))$. By Observation 49, this is at most $(z^{\prime}(\beta)\nu+2\nu^{2})+(-z^{\prime}(\beta)\nu+2\nu^{2})=4\nu^{2}=O(\varepsilon^{2}/n^{2})$.
For (d), we calculate:
$\mu_{0}(0)$ $=\frac{c_{0}}{Z(0)}=\prod_{k=0}^{n-1}\frac{e^{k}}{e^{k}+1}\geq\prod_{k=0}^{n-1}\exp(-e^{-k})\geq\exp(-\sum_{k=0}^{\infty}e^{-k})=e^{-e/(e-1)}\geq 1/5.$
$\mu_{\beta_{\max}}(n)$ $=\frac{c_{n}}{Z(n)}=\prod_{k=0}^{n-1}\frac{e^{n}}{e^{k}+e^{n}}\geq\prod_{k=0}^{n-1}\exp(-e^{-(n-k)})\geq e^{-e/(e-1)}\geq 1/5.\qed$
Proposition 51.
For $n\geq n_{0},q>q_{0},\varepsilon<1/100$, where $n_{0},q_{0}$ are absolute constants, the instance can be constructed so that both ${P^{{\varepsilon}}_{\tt ratio-pt}}$ or ${P^{1/10,\varepsilon}_{\tt count-weak}}$ require cost $\Omega(\frac{\min\{q,n^{2}\}\log\frac{1}{\gamma}}{\varepsilon^{2}})$.
Proof.
By reducing $n$ as needed, and taking $n_{0},q_{0}$ sufficiently large, it suffices to show this under the assumption $10100<n<\sqrt{q}/$ 10. We use Corollary 46 for both problems. Here $\kappa=n\nu=10\varepsilon$, and so $z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}+1/5\leq q$ for $q_{0}$ sufficiently large.
For ${P^{{\varepsilon}}_{\tt ratio-pt}}$ we have $|z(\beta_{\min},\beta_{\max})-z^{(\pm 1)}(\beta_{\min},\beta_{\max})|>2\varepsilon$ as required.
For ${P^{1/10,\varepsilon}_{\tt count-weak}}$ take witnesses $x=0,y=n$. We have $\Delta(x),\Delta(y)\geq 1/5\geq 1/10\cdot e^{2\kappa}$ and we have $|\lambda^{(1)}_{x}-\lambda^{(1)}_{y}|=n\nu=10\varepsilon$ as required.
So both problems require cost $\Omega(\frac{\log(1/\gamma)}{\Psi})=\Omega(\frac{\log(1/\gamma)n^{2}}{\varepsilon^{2}})$. ∎
Combined with Proposition 48, this shows Theorem 47(b,c).
8.2 Bounds for ${P_{\tt count-weak}}$ in the general integer setting
Define $m=\lfloor n/2\rfloor$. We have $d=m$ instances, where we set $\beta_{\min}=0,\beta_{\max}=m\log 2$. We define $c$ by:
$c_{2i}$ $=2^{-i^{2}}$ for $i=0,\dots,m$
$c_{2i-1}$ $=2^{i-i^{2}}\cdot 8\delta$ for $i=1,\dots,m$
For each $i=1,\dots,m$, we form $\lambda^{(i)}$ as
$\lambda^{(i)}_{2i-1}=3\varepsilon,\qquad\qquad\lambda^{(i)}_{x}=0\text{ for all other values $x\neq 2i-1$}$
Let us provide some intuition about this instance. For each $i=1,\dots,m$, there is a small count $c_{2i-1}$ hidden between two larger counts $c_{2i-2},c_{2i}$. The only way to determine the precise value of $c_{2i-1}$ is to take roughly $\frac{1}{\delta\varepsilon^{2}}$ samples of $\mu_{\beta}$ where $\beta\approx i\log 2$. If $\beta\gg i\log 2$, then $\mu_{\beta}(2i)$ will drown out $\mu_{\beta}(2i-1)$, and likewise if $\beta\ll i\log 2$ then $\mu_{\beta}(2i-2)$ will drown out $\mu_{\beta}(2i-1)$. Consequently, to solve ${P^{\delta,\varepsilon}_{\tt count}}$, we need roughly $\Omega(\frac{m}{\delta\varepsilon^{2}})$ samples in total.
Observation 52.
Suppose $\delta<1/100$. Then for any integer $k\in\mathbb{Z}$, we have $Z(k\log 2)\leq 2^{k^{2}+2}$.
Proof.
Here we calculate
$Z(k\log 2)=\sum_{i=0}^{m}2^{-i^{2}}e^{2i(k\log 2)}+8\delta\sum_{i=1}^{m}2^{i-i^{2}}e^{(2i-1)(k\log 2)}\leq\sum_{i=-\infty}^{\infty}2^{-i^{2}+2ik}+8\delta\sum_{i=-\infty}^{\infty}2^{i-i^{2}+(2i-1)k}.$
Terms in the first sum decay at rate at least $1/2$ away from the peak value $i=k$, while the terms in the second sum decay at rate at least $1/4$ from their peak values at $i=k,k+1$. So $Z(\beta)\leq(3\cdot 2^{k^{2}}+8\delta\cdot\frac{8}{3}2^{k^{2}})$, which is less than $2^{k^{2}+2}$ for $\delta<1/100$. ∎
Proposition 53.
Suppose $n\geq 7,\varepsilon<1/100,\delta<1/100$ for large enough constant $n_{0}$. Then we have:
(a) $\Psi\leq O(\delta\varepsilon^{2})$
(b) $z(\beta_{\min},\beta_{\max})\leq n^{2}/4$
(c) $\Delta(0)\geq 1/4$
(d) $\Delta(2k-1)\geq 2\delta$ for all $k=1,\dots,m$.
Proof.
For any $\beta\geq 0$ and $i=1,\dots,m$, we have:
$\frac{Z^{(+i)}(\beta)Z^{(-i)}(\beta)}{Z(\beta)^{2}}=\bigl{(}1+\frac{(e^{3\varepsilon}-1)2^{i-i^{2}}\cdot 8\delta e^{(2i-1)\beta}}{Z(\beta)}\bigr{)}\bigl{(}1+\frac{(e^{-3\varepsilon}-1)2^{i-i^{2}}\cdot 8\delta e^{(2i-1)\beta}}{Z(\beta)}\bigr{)}$
$\qquad\leq\exp\bigl{(}\frac{(e^{3\varepsilon}-1)2^{i-i^{2}}\cdot 8\delta e^{(2i-1)\beta}}{Z(\beta)}\bigr{)}\cdot\exp\bigl{(}\frac{(e^{-3\varepsilon}-1)2^{i-i^{2}}\cdot 8\delta e^{(2i-1)\beta}}{Z(\beta)}\bigr{)}\qquad\text{(since $1+x\leq e^{x}$)}$
$\qquad=\exp\Bigl{(}\frac{8\delta(e^{3\varepsilon}+e^{-3\varepsilon}-2)(2^{i-i^{2}}e^{(2i-1)\beta})}{Z(\beta)}\Bigr{)}$
Since $\varepsilon<1/100$ we have $e^{3\varepsilon}+e^{-3\varepsilon}-2\leq 10\varepsilon^{2}$. Also, $Z(\beta)\geq 2^{-i^{2}}e^{2i\beta}+2^{-(i-1)^{2}}e^{2(i-1)\beta}$ from counts $c_{2i}$ and $c_{2i-2}$. So, we can upper-bound the term $\sum_{r=1}^{d}\log\frac{Z^{(+r)}(\beta)Z^{(-r)}(\beta)}{Z(\beta)^{2}}$ in the definition of $\Psi$ by:
$80\delta\varepsilon^{2}\sum_{i=1}^{m}\frac{2^{i-i^{2}}e^{(2i-1)\beta}}{2^{-i^{2}}e^{2i\beta}+2^{-(i-1)^{2}}e^{2(i-1)\beta}}=80\delta\varepsilon^{2}\sum_{i=0}^{m-1}\frac{2^{i}e^{-\beta}}{1+(2^{i}e^{-\beta})^{2}/2}\leq O(\delta\varepsilon^{2}).$
For (b), Observation 52 gives $Z(\beta_{\max})=Z(m\log 2)\leq 2^{m^{2}+4}$, and trivially we have $Z(\beta_{\min})\geq c_{0}=1$. So $z(\beta_{\min},\beta_{\max})\leq(m^{2}+4)\log 2$, which is smaller than $n^{2}/4$ since $m\leq n/2$ and $n\geq 7$.
For (c), Observation 52 gives $Z(0)\leq 4$. So $\Delta(0)\geq\mu_{0}(0)\geq 1/4$.
For (d), we take $\beta=k\log 2$; Observation 52 gives $\mu_{\beta}(2k-1)=\frac{c_{2k-1}e^{(2k-1)\beta}}{Z(\beta)}\geq\frac{8\delta\cdot 2^{k^{2}}}{2^{k^{2}+2}}=2\delta.$
Proposition 54.
For $n\geq n_{0},q>q_{0},\delta<1/100,\varepsilon<1/100$, where $n_{0},q_{0},\delta_{0}$ are absolute constants, the instance can be constructed so that ${P^{\delta,\varepsilon}_{\tt count-weak}}$ requires cost $\Omega(\frac{\min\{\sqrt{q},n\}\log(1/\gamma)}{\delta\varepsilon^{2}})$.
Proof.
By reducing $n$ as needed, and taking $n_{0},q_{0}$ sufficiently large, it suffices to show this under the assumption that $n<\sqrt{q}/10$. In this case, we apply Corollary 46(b). Here $\kappa=3\varepsilon$. Thus, $z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}/4+6\varepsilon\leq q$ for $q_{0}$ sufficiently large.
For each instance $i$, we use witnesses $x=0,y=2i-1$, where $\Delta(x),\Delta(y)\geq 2\delta\geq\delta e^{2\kappa}$ for $\delta<1/100$, and $|\lambda_{x}^{(i)}-\lambda_{y}^{(i)}|=3\varepsilon$. So ${P^{\delta,\varepsilon}_{\tt count-weak}}$ has cost $\Omega(\frac{d\log(1/\gamma)}{\Psi})\ =\Omega\bigl{(}\frac{n\log(1/\gamma)}{\delta\varepsilon^{2}}\bigr{)}$. ∎
Combined with Theorem 47(b) already shown, this shows Theorem 47(a).
8.3 Bounds for the continuous setting
In the constructions of Sections 8.1 and 8.2 for the integer setting, the complexity bounds have a term $\min\{q,n^{2}\}$. In the continous setting, this dependence on $n$ can be removed; we can get lower bounds which depend only on $q$. Namely, given some family $c^{(i)}$ of integer-setting distributions (supported on $\{0,1,\dots,n\})$, we can construct a new continous-setting instance via:
$\tilde{c}_{x}=c_{n(x-1)},\tilde{\lambda}^{(i)}_{x}=\lambda^{(i)}_{n(x-1)}\text{ for all $x,i$},\qquad\qquad\tilde{\beta}_{\min}=\beta_{\min}n,\tilde{\beta}_{\max}=\beta_{\max}n.$
Crucially here, the new instance $\tilde{c}$ is supported on $x\in[1,2]$, i.e. it has parameter $\tilde{n}=2$. Furthermore observe that
$\tilde{Z}^{(i)}(\beta)=\sum_{x}\tilde{c}_{x}^{(i)}e^{\beta x}=\sum_{x}c^{(i)}_{n(x-1)}e^{\beta x}=\sum_{y}c_{y}^{(i)}e^{\beta(1+y/n)}=e^{\beta}Z^{(i)}(\beta/n).$
So $\tilde{Q}^{(i)}(\tilde{\beta}_{\max})=e^{n(\beta_{\max}-\beta_{\min})}Q(\beta_{\max})$, and $\tilde{\mu}_{\beta n}^{(i)}(1+x/n)=\mu_{\beta}^{(i)}(x)$. Thus, solving ${P^{{\varepsilon}}_{\tt ratio-pt}}$ or ${P^{\delta,\varepsilon}_{\tt count-weak}}$ on the instances $\tilde{c}$ is equivalent to solving them on the original instances $c$. Furthermore, in order to use Corollary 46 on the new problem instances $\tilde{c}$, the original instances $c$ must satisfy
$n(\beta_{\max}-\beta_{\min})+z(\beta_{\min},\beta_{\max})+2\kappa\leq q$
To get the lower bound $\Omega(\frac{q}{\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{{\varepsilon}}_{\tt ratio-pt}}$ and ${P^{\delta,\varepsilon}_{\tt count-weak}}$ in the continuous setting, we construct the problem instances $c$ from Section 8.1 for value $n=\lfloor\sqrt{q}/10\rfloor$. Lemma 50 gives $n(\beta_{\max}-\beta_{\min})+z(\beta_{\min},\beta_{\max})+2\kappa\leq n^{2}+n^{2}+10\varepsilon$, which is at most $q$ for $q\geq q_{0}$ sufficiently large.
Similarly, for the lower bound $\Omega(\frac{\sqrt{q}}{\delta\varepsilon^{2}}\log\tfrac{1}{\gamma})$ for ${P^{\delta,\varepsilon}_{\tt count-weak}}$ in the continous setting, we construct the problem instances $c$ of Section 8.2 for value $n=\lfloor\sqrt{q}/10\rfloor$. For $m=\lfloor n/2\rfloor$, we have $n(\beta_{\max}-\beta_{\min})+z(\beta_{\min},\beta_{\max})+2\kappa\leq n(m\log 2-0)+n^{2}/4+3\varepsilon$, which is at most $q$ for $q\geq q_{0}$ sufficiently large.
This concludes the proof of Theorems 47(d,e).
Appendix A Proofs for error bounds of ${P^{\delta,\varepsilon}_{\tt count}}$
Proof of Theorem 5. Note that $\mu_{\alpha}(x)=\pi(x)e^{\alpha x}/Q(\alpha)$. If $\pi(x)=0$, then the approximation guarantee of ${P^{\delta,\varepsilon}_{\tt count}}$ directly gives $\hat{\mu}_{\alpha}(x)=0=\mu_{\alpha}(x)$. Otherwise, we can use the approximation guarantees of ${P^{\delta,\varepsilon}_{\tt count}}$ and ${P^{{\varepsilon}}_{\tt ratio-all}}$ to show the upper bound on $\hat{\mu}_{\alpha}(x)$:
$\hat{\mu}_{\alpha}(x)$ $\leq\frac{e^{\varepsilon}\pi(x)(1+\varepsilon(1+\delta/\Delta(x)))e^{\alpha x}}{Q(\alpha)}=e^{\varepsilon}\mu_{\alpha}(x)(1+\varepsilon(1+\delta/\Delta(x)))$
$\leq e^{\varepsilon}\mu_{\alpha}(x)(1+\varepsilon(1+\delta/\mu_{\alpha}(x)))\qquad\qquad\qquad\qquad\qquad\qquad(\text{$\mu_{\alpha}(x)\leq\Delta(x)$ by definition})$
$=e^{\varepsilon}(1+\varepsilon)\mu_{\alpha}(x)+e^{\varepsilon}\delta\leq\mu_{\alpha}(x)+3\varepsilon(\mu_{\alpha}(x)+\delta)\qquad\qquad\ (\text{straightforward calculus for $\varepsilon<1/2$})$
The lower bound is completely analogous.
Proof of Theorem 6. Let us show $\hat{Q}(\alpha\mid\mathcal{D})\leq e^{\varepsilon}Q(\alpha)$ for all $\alpha$; the lower bound $\hat{Q}(\alpha\mid\mathcal{D})\geq e^{-\varepsilon}Q(\alpha)$ is completely analogous. For each value $i$ with $c_{i}\neq 0$, the guarantee of problem ${P^{1/n,\varepsilon/3}_{\tt count}}$ gives $\hat{\pi}(i)\leq\pi(i)+\varepsilon/3\cdot\pi(i)\bigl{(}1+\frac{1}{n\Delta(i)}\bigr{)}$. Since also $\hat{\pi}(i)=\pi(i)$ when $c_{i}=0$, we can calculate
$\hat{Q}(\alpha\mid\mathcal{D})\leq\sum_{i}\pi(i)(1+\varepsilon/3)e^{\alpha i}+\varepsilon/3\sum_{i:c_{i}\neq 0}\frac{\pi(i)e^{\alpha i}}{n\Delta(i)}$ (26)
Note that $Q(\alpha)=\sum_{i}\pi(i)e^{\alpha i}$. Thus, the first term in the RHS of (26) is $(1+\varepsilon/3)Q(\alpha)$. We also observe that $\Delta(i)\geq\mu_{\alpha}(i)=\frac{\pi(i)e^{\alpha i}}{Q(\alpha)}$, so $\sum_{i:c_{i}\neq 0}\frac{\pi(i)e^{\alpha i}}{n\Delta(i)}\leq\sum_{i:c_{i}\neq 0}\frac{Q(\alpha)}{n}\leq Q(\alpha)$. Overall we have $\hat{Q}(\alpha\mid\mathcal{D})\leq(1+\varepsilon/3)Q(\alpha)+\varepsilon/3\cdot Q(\alpha)\leq e^{\varepsilon}Q(\alpha)$ as desired.
Appendix B Correctness with approximate oracles
Here, we define the total variation distance $||\cdot||_{TV}$ by
$||\tilde{\mu}_{\beta}-\mu_{\beta}||_{TV}=\max_{{\cal K}\subseteq\mathcal{F}}|\tilde{\mu}_{\beta}({\cal K})-\mu_{\beta}({\cal K})|=\tfrac{1}{2}\sum_{x}|\tilde{\mu}_{\beta}(x)-\mu_{\beta}(x)|$
Theorem 55.
Suppose that sampling procedure $\mathfrak{A}$ has cost at most $T$ for some known parameter $T$. Let $\tilde{\mathfrak{A}}$ be the algorithm obtained from $\mathfrak{A}$ as follows: (i) we replace calls $x\sim\mu_{\beta}$ with calls $x\sim\tilde{\mu}_{\beta}$ where $||\tilde{\mu}_{\beta}-\mu_{\beta}||_{TV}\leq\gamma/T$; (ii) we terminate algorithm after $T/\gamma$ steps and return arbitrary answer.
Then $\tilde{\mathfrak{A}}$ has cost $O(T)$, and the total variation distance of the of $\mathfrak{A}$ and $\mathfrak{A}^{\prime}$ is at most $2\gamma$.
Proof.
The distributions $\mu_{\beta}$ and $\tilde{\mu}_{\beta}$ can be maximally coupled so that samples $x\sim\mu_{\beta}$ and $x\sim\tilde{\mu}_{\beta}$ are identical with probability at least $1-\gamma/T$. Assume the $k^{\text{th}}$ call to $\mu_{\beta}$ in $\mathfrak{A}$ is coupled with the $k^{\text{th}}$ call to $\tilde{\mu}_{\tilde{\beta}}$ in $\tilde{\mathfrak{A}}$ when both calls are defined and $\beta=\tilde{\beta}$. We say the $k^{\text{th}}$ call is good if either (i) both calls are defined and the produced samples are identical, or (ii) $\mathfrak{A}$ has terminated earlier. Note that ${\mathbb{P}}[\mbox{$k^{\text{th}}$ call is good}\>|\>\mbox{all previous calls were good}]\geq 1-\gamma/T$, since the conditioning event implies $\beta=\tilde{\beta}$ (assuming the calls are defined).
Let $A$ and $\tilde{A}$ be the number of calls to the sampling oracle by algorithms $\mathfrak{A}$ and $\tilde{\mathfrak{A}}$, respectively. By assumption, ${\mathbb{E}}[A]=T$. We say that the execution is good if three events hold:
  1. $\mathcal{E}_{1}$:
[…]
  2. $\mathcal{E}_{2}$:
    The number of oracle calls by $\mathfrak{A}$ is at most $T/\gamma$. By Markov’s inequality, this has probability at least $1-\frac{{\mathbb{E}}[A]}{T/\gamma}\geq 1-\gamma$.
If these two events occur, then $\tilde{\mathfrak{A}}$ returns the same output as $\mathfrak{A}$; by the union bound, this has probability at least $1-2\gamma$. Thus, $\tilde{\mathfrak{A}}$ and $\mathfrak{A}$ have total variation distance at most $2\gamma$.
[…]
Appendix C Proofs for statistical sampling
Proof of Lemma 13. When $p\geq\bar{p}$ we have $F(Np,N\varepsilon(p+\bar{p}))\leq F(Np,N\varepsilon p)\leq 2e^{-Np\varepsilon^{2}/3}$; for $N$ larger than the stated bound and $p\geq\bar{p}$ this is at most $\gamma/2$. When $p<\bar{p}$, we have $F(Np,N\varepsilon(p+\bar{p}))\leq F(Np,N\varepsilon\bar{p})$; by monotonity this is at most $F(N\bar{p},N\varepsilon\bar{p})\leq\gamma/2$.
For the second bound, we use standard estimates $F_{+}(Np,Nx)\leq\exp(\frac{-Nx^{2}}{2(p+x)})$ with $x=p(e^{\varepsilon}-1)$ and $F_{-}(Np,Nx)\leq\exp(\frac{-Nx^{2}}{2p})$ with $x=p(1-e^{-\varepsilon})$. These show that, for $N$ larger than the stated bound and $p\geq e^{-\varepsilon}\bar{p}$, the upper and lower deviations have probability at most $\gamma/4$. When $p\leq e^{-\varepsilon}\bar{p}$, we use the monotonicity property $F_{+}(Np,N(\bar{p}-p))\leq F_{+}(Ne^{\varepsilon}\bar{p},N(\bar{p}-e^{-\varepsilon}\bar{p}))$.
Proof of Lemma 14. For brevity, let $a=\mu_{\alpha}(x),\hat{a}=\hat{\mu}_{\alpha}(x),\eta=Q(\alpha)e^{-\alpha x},\hat{\eta}=\hat{Q}(\alpha)e^{-\alpha x}$ and $y=\delta/\Delta(x)$. If $a=0$, then $\hat{\pi}(x)=0$ as desired. So suppose $a>0$. For the upper bound, we calculate
$\frac{\hat{\pi}(x)-\pi(x)}{\pi(x)}$ $=\frac{\hat{\eta}\hat{a}}{\eta a}-1\leq\frac{\eta e^{\varepsilon/3}\cdot(a+\varepsilon a/3\cdot(1+y))}{\eta a}-1$ (by (A1) and (A2))
$=\bigl{(}e^{\varepsilon/3}\cdot(1+\varepsilon/3)-1\bigr{)}+\bigl{(}1/3\cdot e^{\varepsilon/3}\bigr{)}\varepsilon y$
$\leq\varepsilon+\varepsilon y=\varepsilon(1+\delta/\Delta(x))$ (straightforward calculus)
The complementary inequality is completely analogous.
Proof of Theorem 15. Consider the following procedure: for each $i=1,\dots,N$ draw $r=\lceil 100\tau\rceil$ copies of $X_{i}$ and compute the sample average $\overline{X_{i}}$. For any pair $i^{\prime}\leq i$, define $Y_{i^{\prime},i}=\prod_{\ell=i^{\prime}}^{i}\overline{X_{\ell}}$.
We can first observe that, for any values $i,i^{\prime}$, we have ${{\mathbb{P}}\!\bm{\left[}Y_{i^{\prime},i}/{\mathbb{E}}[Y_{i^{\prime},i}]\in[e^{-\varepsilon/2},e^{\varepsilon/2}]\bm{}}\geq 0.93$. For, let $Y=Y_{i^{\prime},i}$ for brevity. Since multiplying the variables $X_{j}$ by constants does not affect the claim, we may assume that ${\mathbb{E}}[X_{j}]=1$. Each variable $\overline{X_{j}}$ is the mean of $r$ independent copies of $X_{j}$, so ${\mathbb{E}}[\overline{X_{j}}]={\mathbb{E}}[X_{j}]=1$ and $\text{Variance}[\overline{X_{j}}]=\text{Variance}[X_{j}]/r={{\mathbb{S}}[{X_{j}}]}/r$. So ${\mathbb{E}}[Y]=1$, and we have:
${\mathbb{E}}[Y^{2}]=\prod_{\ell=i^{\prime}}^{i}(1+\text{Variance}[\overline{X_{\ell}}])=\prod_{\ell=i^{\prime}}^{i}(1+\frac{\text{Variance}[X_{\ell}]}{r})\leq e^{\sum_{\ell=i^{\prime}}^{i}{{\mathbb{S}}[{X_{\ell}}]}/r}\leq e^{(\tau\varepsilon^{2})/r}\leq e^{\varepsilon^{2}/100}$
It now suffices to show $|Y-1|\leq(1-e^{-\varepsilon/2})$. By Chebyshev’s inequality, ${{\mathbb{P}}\!\bm{\left[}|Y-1|>(1-e^{-\varepsilon/2})\bm{}}\leq\frac{\text{Variance}[Y]}{(1-e^{-\varepsilon/2})^{2}}\leq\frac{e^{\varepsilon^{-2}/100}-1}{(1-e^{-\varepsilon/2})^{2}}$. Straightforward calculations show that this is at most $0.07$ for $\varepsilon\in(0,1)$.
[…]
We can execute this procedure for $k=O(\log\tfrac{1}{\gamma})$ independent trials and return
$\hat{X}^{\tt prod}_{i}=\text{median}(Y_{1,i}^{(1)},\dots,Y_{1,i}^{(k)})\qquad\text{for $i=1,\dots,N$}$
where $Y_{1,i}^{(j)}$ is the value of statistic $Y_{1,i}$ in the $j^{\text{th}}$ trial. With probability at least $1-\gamma$ the bad-event $\mathcal{E}^{(j)}$ holds in fewer than $k/2$ trials $j$. In this case $\hat{X}^{\tt prod}$ satisfies the required condition for all $i$.
Appendix D The ${\tt Balance}$ subroutines
The starting point for the algorithm is a sampling procedure of Karp & Kleinberg [20] for noisy binary search, which we summarize as follows: [^5] The original algorithm in [20] only gives success probability $3/4$; it is well-known that it can be amplified to $1-\gamma$ by checking the returned solution and restarting as needed.
Theorem 56 ([20]).
Suppose we can sample from Bernoulli random variables $X_{0},\dots,X_{N}$, wherein each $X_{i}$ has some unknown mean $x_{i}$, and $0\leq x_{0}\leq x_{1}\leq x_{2}\leq\dots\leq x_{N}\leq 1$. Let us also write $x_{-1}=0,x_{N+1}=1$. For any $\alpha,\nu,\gamma\in(0,1)$, there is a procedure which uses $O(\frac{\log(N/\gamma)}{\nu^{2}})$ samples from the variables $X_{i}$ in expectation. With probability at least $1-\gamma$, it returns an index $V\in\{-1,0,1,\dots,N\}$ such that $x_{V}-\nu\leq\alpha\leq x_{V+1}+\nu$.
Consider the non-decreasing function $p(\beta)=\mu_{\beta}[\chi,n]=\frac{\sum_{x\geq\chi}c_{x}e^{\beta x}}{Z(\beta)}$. Our basic plan is to quantize the search space $[\beta_{\min},\beta_{\max}]$ into discrete points $u_{0},\dots,u_{N}$, for some value $N=\operatorname*{poly}(n,q)$, and then apply Theorem 56 for probabilities $x_{i}=p(u_{i})$. Note that we can simulate a Bernoulli random variable of rate $x_{i}$ by drawing $y\sim\mu_{u_{i}}$ and checking if $y\geq\chi$. The full details are as follows:
Procedure ${\tt Balance}(\chi,\tau,\gamma)$.
1 if $\chi\leq 0$ then return $\beta=\beta_{\min}$
2 Set $\beta_{\min}^{\prime}=\max\{\beta_{\min},\beta_{\max}-(q+1)\}$ and $N=\Bigl{\lceil}\frac{n(q+1)}{2\log\bigl{(}\frac{4-4\tau}{3-2\tau}\bigr{)}}\Bigr{\rceil}$
3 for $j=0,\dots,N$ set $u_{j}=(1-j/N)\beta_{\min}^{\prime}+(j/N)\beta_{\max}$
4 Apply Theorem 56 for values $x_{j}=p(u_{j})$ with $\alpha=\tfrac{1}{2},\nu=(\tfrac{1}{2}-\tau)/2$; let $V$ be the return value.
5 if $-1<V<N$ then return $\beta=\frac{u_{V}+u_{V+1}}{2}$
6 else if $V=-1$ then return $\beta=\beta_{\min}^{\prime}$
else if $V=N$  then return $\beta=\beta_{\max}$
This algorithm has complexity $O(\log\frac{N}{\gamma})=O(\log\frac{nq}{\gamma})$ for fixed $\tau$. It is clearly correct if $\chi\leq 0$ since then $\mu_{\beta_{\min}}[\chi,n]=1$. For $\chi>0$, we claim that if $V$ satisfies $x_{V}-\nu\leq\alpha\leq x_{V+1}+\nu$, which occurs with probability at least $1-\gamma$, then $\beta\in{\tt BalancingVals}(\chi,\tau)$. There are a number of cases.
  • Suppose that $-1<V<N$. It suffices to show that $\tau\leq p(\beta)\leq 1-\tau$. We will show only the inequality $p(\beta)\leq 1-\tau$; the complementary inequality is completely analogous.
    The condition $x_{V}-\nu\leq\alpha$ implies that $p(u_{V})\leq 1/2+\nu$. We have $\beta-u_{V}=\frac{\beta_{\max}-\beta_{\min}^{\prime}}{2N}\leq\frac{\log\bigl{(}\frac{4-4\tau}{3-2\tau}\bigr{)}}{n}$, and so we get
    $p(\beta)=p(u_{V})\cdot\frac{Z(u_{V})}{Z(\beta)}\cdot\frac{\sum_{x\in[\chi,n]}c_{x}e^{\beta x}}{\sum_{x\in[\chi,n]}c_{x}e^{u_{V}x}}\leq p(u_{V})\cdot 1\cdot e^{(\beta-u_{V})n}\leq(1/2+\nu)\cdot\frac{4-4\tau}{3-2\tau}=1-\tau.$
  • Suppose that $V=-1$. Here, the condition $x_{V+1}+\nu\geq\alpha$ implies that $p(u_{0})\geq 1/2-\nu$. So for the lower bound on $p(\beta)$, we have $p(\beta)=p(u_{0})\geq 1/2-\nu\geq\tau$.
    We have either $\beta=\beta_{\min}$ or $\beta=\beta_{\max}-(q+1)$. In the former case, we do not need to show the upper bound on $p(\beta)$. In the latter case, we have
    $p(\beta)=p(\beta_{\max})\cdot\frac{Z(\beta_{\max})}{Z(\beta)}\cdot\frac{\sum_{x\in[\chi,n]}c_{x}e^{\beta x}}{\sum_{x\in[\chi,n]}c_{x}e^{\beta_{\max}x}}\leq 1\cdot\frac{Z(\beta_{\max})}{Z(\beta_{\min})}\cdot e^{-(\beta_{\max}-\beta)}$
    where the last inequality holds since $c_{x}=0$ for $x\in(0,1)$ and since $\chi>0$. Since $\beta=\beta_{\max}-(q+1)$, we conclude that $p(\beta)\leq\frac{Z(\beta_{\max})}{Z(\beta_{\min})}\cdot e^{-(q+1)}\leq e^{-1}\leq 1/2$.
  • Suppose that $V=N$. Since $\beta=\beta_{\max}$, we only need to show the upper bound on $p(\beta)$. Here, the condition $x_{V}-\nu\leq\alpha$ implies $p(\beta)=p(u_{V})\leq 1/2+\nu\leq 1-\tau$.
This concludes the proof of Theorem 11. We next describe the accelerated algorithm given access to a data structure for ${P^{{\varepsilon}}_{\tt ratio-all}}$ where $\varepsilon=\frac{1/2-\tau}{30}$. Here, the search space is reduced to just $N=O(q)$.
Procedure ${\tt BalancePreprocessed}(\chi,\tau,\mathcal{D},\gamma)$ 1 if $\log\hat{Q}(\beta_{\max}\mid\mathcal{D})>q+\varepsilon$ then terminate with ERROR. 2 set parameters $N=\Bigl{\lfloor}\frac{\log\hat{Q}(\beta_{\max}\mid\mathcal{D})}{10\varepsilon}\Bigr{\rfloor}$, and set $u_{0}=\beta_{\min},u_{N}=\beta_{\max}$ 3 for $j=1,\dots,N-1$ let $u_{j}\in[\beta_{\min},\beta_{\max}]$ be an arbitrary value with $|\log\hat{Q}(u_{j}\mid\mathcal{D})-10\varepsilon j|\leq\varepsilon$                                                   (if no such value exists then terminate with ERROR). 4 Apply Theorem 56 for values $x_{j}=p(u_{j})$ with $\alpha=\tfrac{1}{2},\nu=(\tfrac{1}{2}-\tau)/2$; let $V$ be the return value. 5 if $-1<V<N$ then return $\beta=u_{V+1}$ 6 else if $V=-1$ then return $\beta=\beta_{\min}$ else if $V=N$  then return $\beta=\beta_{\max}$
Due to the check at Line 1, the algorithm has complexity $O(\log\frac{q}{\gamma})$ as desired. For correctness, we suppose that $\mathcal{D}$ is good. In this case, the check at Line 2 will not terminate with error; also, for each $j=1,\dots,N-1$ the value $u$ with $Q(u)=e^{10\varepsilon j}$ will work at Line 4; hence this check does not fail either. Furthermore, for every value $j=0,\dots,N-1$ we have $|\log Q(u_{j})-10\varepsilon j|\leq 2\varepsilon$ and similarly $\log Q(u_{N})-\log Q(u_{N-1})\in[8\varepsilon,22\varepsilon]$. So, for all $j=0,\dots,N$, we have
$z(u_{j},u_{j+1})\in[6\varepsilon,22\varepsilon]$ (27)
So indeed $0\leq x_{0}\leq\dots\leq x_{N}\leq 1$ as required. To finish the proof, we claim that if $V$ satisfies $x_{V}-\nu\leq\alpha\leq x_{V+1}+\nu$, which occurs with probability at least $1-\gamma$, then $\beta\in{\tt BalancingVals}(\chi,\tau)$.
  • Suppose that $-1<V<N$. The inequality $x_{V+1}+\nu\geq\alpha$ implies that $p(\beta)=p(u_{V+1})\geq\alpha-\nu\geq\tau$. The inequality $x_{V}-\nu\leq\alpha$ implies that $p(u_{V})\leq\alpha+\nu$. From Eq. (27), we can calculate
    $1-p(\beta)=(1-p(u_{V}))\cdot\frac{Z(u_{V})}{Z(u_{V+1})}\cdot\frac{\sum_{x<\chi}c_{x}e^{u_{V+1}x}}{\sum_{x<\chi}c_{x}e^{u_{V}x}}\geq(1-\alpha-\nu)\cdot e^{-22\varepsilon}\cdot 1$
    and straightforward calculations show that this is at least $\tau$.
  • Suppose that $V=-1$. The condition $x_{V+1}+\nu\geq\alpha$ implies that $p(\beta)=p(u_{0})\geq 1/2-\nu\geq\tau$. Since $\beta=\beta_{\min}$, we do not need to show the upper bound on $p(\beta)$.
  • Suppose that $V=N$. The condition $x_{V}-\nu\leq\alpha$ implies that $p(\beta)=p(u_{N})\leq 1/2+\nu\leq 1-\tau$. Since $\beta=\beta_{\max}$, we do not need to show the lower bound on $p(\beta)$.
This concludes the proof of Theorem 11.
Appendix E Some useful inequalities
Lemma 57.
Let $b_{1},\ldots,b_{m+1}$ be an strictly increasing positive sequence. Then we have $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}\geq\frac{m^{2}}{\log(b_{m+1}/b_{1})}-m/2$.
Proof.
Define $v=\log(b_{m+1}/b_{1})$. For $i=1,\dots,m$ let $a_{i}=\log(b_{i+1}/b_{i})$. Then $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}=\sum_{i=1}^{m}\frac{1}{e^{a_{i}}-1}$. By telescoping sums we have $\sum_{i}a_{i}=\log(b_{2}/b_{1})+\log(b_{3}/b_{2})+\dots+\log(b_{m+1}/b_{m})=v$.
By Jensen’s inequality applied to the concave-up function $x\mapsto\frac{1}{e^{x}-1}$, we have
$\sum_{i=1}^{m}\frac{1}{e^{a_{i}}-1}\geq m\cdot\frac{1}{e^{\tfrac{1}{m}\sum_{i=1}^{m}a_{i}}-1}=\frac{m}{e^{v/m}-1}.$
By straightforward calculus, it can be verified that $\frac{1}{e^{y}-1}\geq\frac{1}{y}-\frac{1}{2}$ for all $y\geq 0$. Applied to $y=v/m$, we conclude that $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}\geq m\bigl{(}m/v-1/2\bigr{)}=m^{2}/v-m/2$ as claimed. ∎
Lemma 58.
Let $b_{1},\ldots,b_{m}$ be a log-concave sequence with $b_{k}\leq\frac{1}{k}$ for each $k$. Then $b_{1}+\ldots+b_{m}<e$.
Proof.
Let $k\in\{1,\dots,m\}$ be chosen to maximize the value $kb_{k}$ (breaking ties arbitrarily). If $b_{k}=0$, then due to maximality of $k$ we have $b_{1}=\dots=b_{m}=0$ and the result holds. Suppose that $1<k<m$; the cases with $k=1$ and $k=m$ are very similar. Define the sequence $y_{1},\dots,y_{m}$ by:
$y_{i}=\begin{cases}\frac{1}{k}(\frac{k-1}{k})^{i-k}&\mbox{if }i<k\\ \frac{1}{k}(\frac{k}{k+1})^{i-k}&\mbox{if }i\geq k\end{cases}$
By maximality of $k$, we have $\frac{y_{k-1}}{y_{k}}=\frac{k}{k-1}\geq\frac{b_{k-1}}{b_{k}}$ and $\frac{y_{k+1}}{y_{k}}=\frac{k}{k+1}\geq\frac{b_{k+1}}{b_{k}}$ and $y_{k}=\frac{1}{k}\geq b_{k}$. Since $\log y_{i}$ is linear for $i\leq k$ and for $i\geq k$, log-concavity of sequence $b$ gives $b_{i}\leq y_{i}$ for $i=1,\dots,m$. So
$\sum_{i=1}^{m}b_{i}$ $\leq\sum_{i=1}^{\infty}y_{i}=\sum_{i=1}^{k-1}\frac{1}{k}\Bigl{(}\frac{k-1}{k}\Bigr{)}^{i-k}+\sum_{i=k}^{\infty}\frac{1}{k}\Bigl{(}\frac{k}{k+1}\Bigr{)}^{i-k}=\Bigl{(}1-\frac{1}{k}\Bigr{)}^{1-k}+\frac{1}{k}$
Now consider the function $g(x)=(1-x)^{1-1/x}+x$. We have shown $\sum_{i}b_{i}\leq g(1/k)$. To finish the proof, it suffices to show $g(x)<e$ for all $x\in(0,1/2)$. This follows from some standard analysis showing that $\lim_{x\rightarrow 0}g(x)=e$ and $\lim_{x\rightarrow 0}g^{\prime}(x)=1-e/2<0$ and $g^{\prime\prime}(x)<0$ in the interval $(0,\tfrac{1}{2})$. ∎
Acknowledgments
We thank Heng Guo for helpful explanations of algorithms for sampling connected subgraphs and matchings, Maksym Serbyn for bringing to our attention the Wang-Landau algorithm and its use in physics. The author Vladimir Kolmogorov was supported by the European Research Council under the European Unions Seventh Framework Programme (FP7/2007-2013)/ERC grant agreement no 616160.